by TTC TTC

800 – Fire Pump Study

10/19/19 – SOSS received and ThermalTech onsite for fire pump testing. We witnessed the fire pump test by SA Comunale and they stated that the fire pumps passed with both city water mains open.
10/30/19 – Terry and Cai met to review the model and report. So far the model is working and matching existing conditions. Cai did a great job interpolating the existing pressure drops from the field data we collected. It is our initial view that both city water mains need to be open for the fire pumps because of the large pressure drop across the BFP’s. WE have to get the mfgr data to confirm this. We also are setting up additional testing with CDFP for more field test.
12/12/19 – No new updates this month.
01/24/20 – This project can be archived in lieu of the City Water Main replacement which has been submitted for Bidding & Construction.

by THP THP

800 – City Facade Assessment (THP F105)

12/28/17 – THP publish SOSS to Hamilton County as requested.
01/25/18 – Hamilton County approved SOSS.
03/06/2018 – THP completed ground survey of façade.
03/29/2018 – THP to schedule rappelling survey early April.
04/26/18 – THP schedule rappel the facade.
05/21/18 – THP completed draft of Cincinnati façade ordinance report and published.
5/31/18 – County reviewing report.
6/26/18 – THP submitted drafts of the Envelope Assessment Report and City Report
07/26/18 – THP finalized and published Cincinnati façade ordinance report. Signature by Ralph Linne’s required before THP submits report to City of Cincinnati Buildings & Inspections.
08/24/18 – THP received letter from City of Cincinnati acknowledging the submitted report.

by TEI TEI

800 LEED EB+OM Certification

Project Lead: GBBN

05/25/10 – Bert Watts requested SOSS to be prepared. This is second SOSS for overall project. This second SOSS is to await results of Phase 1 Planning Study SOSS to determine complete scope of work for project.

05/26/10 – GBBN requested estimate for consulting services from ThermalTech.

06/24/10 – SOSS submitted to Bert Watts at monthly Partner Meeting.

04/28/11 – SOSS reconfiguration discussed at monthly partnering meeting.

01/25/12 – County and partners completed a planning meeting and established project goals and schedule. Project recording period to end October 31, 2012. Three month recording period August 1 to October 31. County review period for credits June 1 to July 3Design team project meeting schedule for May 10th with County review period of credits starting June 1. THP reviewing credits for use of park across of street from the building. Design team project meeting schedule for May 10th with County review period of credits starting June 1. THP submitted policy’s on June 4th for County review. County in the process of reviewing credits.

05/10/12 – Design Team Status Meeting at GBBN.

05/11/12 – Bert Watts responds to Design Team questions developed in Design Team Meeting.

05/26/11 – GBBN/TEI/THP to discuss options for reconfiguring the scope and services for each firm to perform the Certification process.

06/09/11 – Final SOSS due from each firm. GBBN delivers SOSS F017A Revised to Bert Watts.

07/11/11 – SOSS received approved from Bert Watts.

07/13/11 – Confirmed all three Partners have received approved SOSS.

08/32/11 – Kick-off Meeting to be held with THP and ThermalTech at GBBN offices.

08/03/11 – All partners held a kick-off meeting to discuss logistics and timetables.

09/22/11 – The design team, minus THP, and Bert Watts meet to discuss project schedule generated by TEI and project registration. A meeting with THP is to be held in future.

09/28/11 – Tony Yunker sends invitation to project team members to register on the USGBC Online site.

05/24/12 – Tony Yunker completes preliminary new County Procedures for Owner Review. THP has also completed the sample Procedure Documents.

06/06/12 – TEI discusses refrigerant conversion strategy with Ralph Linne.

06/13/12 – Conference call with all three partners to discuss Procedure strategy for delivery to Facilities Department. THP forwards Murray Monroe summaries and procedures. John Millar yet to complete procedures. Tony Yunker updates spread sheet with performance period.

06/22/12 – County notifies Murray Monroe of their lack of desire to converting parking spaces at east or rear of building to green space for SS Credit 5.

06/25/12 – ThermalTech completed its Preliminary Reviews for County approval. ThermalTech indicates IEQc1.3 is to be removed from possible point list.

09/26/12 – GBBN/THP/TEI hold conference call to discuss delays in receiving procedure review comments back from owner.

09/27/12 – Bert Watts leaves voice mail concerning results of review meeting with Ralph Linne. Bert is to review existing procedure templates for overlaps. Solid Waste Stream Audit may need to be revisited.

01/27/13 – GBBN/THP/TEI agree to establish the final three month credit performance period for May 30, 2013 to August 31,2013. This information passed along to Bert Watts that agreed to these dates.

04/22/13 – County requested THP coordinate Bayer Becker surveyor .

04/05/13 – GBBN, TEI and THP hold design team meeting to discuss scheduling and replacement of Tony Yunker as Administrator.

04/24/13 – GBBN held an internal meeting to get staff onboard with the project in preparation for the County to deliver the required procedure updates. The County is to deliver the updated procedures by the first week of May.

04/26/13 – In4 helped to calculate the building occupant load based on the CAFM data in Archibus.

05/07/13 – GBBN contacts Darnell Edwards, Building Manager, to inform him of the impending presence of Bayer Becker surveyors and their need to open doors for elevation shots of the landings inside the building.

05/30/13 – Bayer Becker completed survey and anticipated providing drawings by the end of the week. ThermalTech finished a couple more water surveys. The building probably qualifies for the higher water conservation level point.

06/06/13 – THP forwards an electronic copy of the final survey by BBE.

06/06/13 – THP issued Bayer Becker survey.

06/13/13 – GBBN inquires about County delays in transmitting revised procedures to Design Team. More revisions are necessary and the Director and Assistant Director must review the proposed changes before transmitting them.

08/06/13 – Design Team met to discuss Procedures returned from Owner and input requirements.

08/07/13 – Emails sent to Darnell Edwards to request meeting to discuss Procedure implementation and forms necessary.

08/26/13 – Bert Watts and Tony Matre confirm Darnell has been on vacation henceforth no response to GBBN emails.

08/28/13 – TTC started the Building Report. TTC and Amy Crousey will be inputting data and forms this week and next for the Sept 5 meeting at GBBN.

09/05/13 – GBBN cancels design team meeting due to inability to contact owner to arrange owner meeting.

09/06/13 – GBBN receives request from Facilities concerning possible meeting date with Building Manager. Meeting is set for 09/12/13.

09/12/13 – GBBN meets with Facilities and Building Manager to discuss implementation issues and necessity for compliance documentation forms.

09/26/13 – TEI input several credits this month in the EA and WE divisions. The County provided some additional PM information and TTC uploaded to update the credit.

10/15/13 – TEI uploaded schedule for PM.

11/15/13 – TEI Updated Credits for water.

12/15/13 – TEI updated credits for EA, IAQ.

01/30/14 – Bert asked for a definitive answer about final submission to USGBC by 1/31 plus an explanation about not meeting the 12/31 deadline.

02/24/14 – ThermalTech entered most WE credits and EA credits. County provided and approved paperwork in these credits for upload. A&M on site creating trends for use with IAQ credits and commissioning credits.

02/26/14 – GBBN sent revised 11.0 Sustainability Purchasing- Facility Alterations Tracking Form back to Bert Watts. Master completion matrix updated, we can uses at partner meeting to track status. Performance Period should be finalized from about June 2012 through either February or March 2014.

03/14/14 – GBBN transfers floor plans, site boundary plan and project photographs to other partners for uploading into the project website.

03/27/14 – Shift Preliminary Submission date to April 7, 2014. Setup meeting with Ralph and Bert.

04/16/14 – Owner Signatory Meeting with Ralph Linne and Bert Watts. Credits and Pre-Requisites requiring owner approvals/signatures were reviewed and modified.

04/21/14 – Bert Watts forwards Consumable and Light Bulb materials to GBBN.

04/24/14 – THP found some more owner checklist items and they will update these after milestone goes without response. Reported to Tony Matre via telephone call.

07/24/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/05/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

08/27/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

02/24/14 – GBBN and Pinnacle Engineering met with Tony Matre to deliver the final reports and electronic copies of the reports. GBBN is completing final report drawings for Pinnacle as final preliminary drawings are delivered by Pinnacle to GBBN. All drawings are to be completed by 2/28/14.

03/26/14 – All asbestos reports and associated drawings are complete. Final drawing input into Archibus/WebCentral is ongoing. GBBN was granted access to the database to complete the drawing downloads.

07/01/14 – GBBN sends another email and follow-up voice mail to Pinnacle Engineering inquiring about status of billings and final completion.

08/28/14 – TEI updated two credit points so far. We have edited about four others. The OA prerequisite is the major one we are working on this week, GBBN has finished updating all the rebuttal questions, THP has some questions about the green cleaning policy for maintenance log.

09/05/14 – TEI updated all WE credits this week. TEI Update all EA credits this week. TEI is working on IEQ credits and outdoor air requirements..

09/12/14 – All partners are updated with LEED input except TEI is working on IEQ input which may require an air balance report on the exhaust fan system. TTC requested time with Darnell to do this work.

09/19/14 – TEI updated all credits but IEQp1 which doesn’t meet minimum standards yet. Final submission date reset to 10/1/14. More work needed in building and programming changes required on AHU’s to meet this prereq.

09/24/14 – Refer to GBBN updated.

09/25/14 – ThermalTech completed the remaining points for USGBC response including all fieldwork for IEQp1 credits yesterday. The tables have been updated. We are now updating the credit and uploading our final forms for final submission on 10/1/2014.

10/22/24 – Two clarifications came back from USGBC in regards to the IEQp1 and EAp2 prerequisites. These are both on the ThermalTech realm to answer. TTC posted two response questions to the website asking for clarification. ThermalTech is currently uploading the 800 Energy Star forms for 2014 for the 88 score. Answers are due back to USGBC by 11/5/14 on these clarifications. ThermalTech should have both answers uploaded prior to that. It is possible to submit EAp2 as a “No Recent Labels” but that seems counterintuitive to what we are trying to accomplish at 800. Our score is up 2 points, we are hopeful the will wait until the certification goes through to give us our final certification.

09/24/14 – Refer to GBBN updated.

10/23/14 – LEED team receives Mid-Review Clarification letter via email. TEI indicates both requests related to their credits.

10/24/14 – TEI sends response back to USGBC.

11/05/14 – Mid Level Review responses due to USGBC for EAp2 and IEQp1. ThermalTech request extension after 3 clarifications for each prereq are still in the process of revision with USGBC.

11/19/14 – TEI sends Final Mid-Level responses back to USGBC concerning prereq credits EAp2 and IEQp1. After about one weeks worth of work at ThermalTech the revised prereq clarifications were sent back to USGBC.

11/28/14 – The LEED Team recommended the County accept the Certified level offered by USGBC given the cost, work, time and efforts involved in any Formal Inquiries about additional points that my be available. The chances are low that this building could reach Silver Level without significant capital cost involved in additional upgrades. HamCo accepts Certified Level, GBBN reports this decision to USGBC and the project is formally Certified Level at 43 points. Final Scorecard: Attempted | Awarded | Final Report. Congratulations to the entire team!!!

by TEI TEI

800 Asbestos Study

Project Lead: GBBN

10/20/11 – Kick-off Meeting with Pinnacle and Facilities Department

11/04/11 – Facilities Department sent a list of buildings to be included in Asbestos Abatement Reporting.

04/04/12 – Pinnacle Engineering requested drawing files for the purpose of documenting findings.

04/05/12 – GBBN sent drawing files to Pinnacle Engineering in PDF form.

03/30/12 – Meeting with Tim Schue of Facilities Department to discuss Archibus Input Strategy

04/05/12 – Meeting with GBBN and Facilities to discuss drawing format.

04/06/12 – Meeting with Building Manager and Pinnacle to kick-off Field Investigation.

05/10/12 – Meeting with GBBN and Pinnacle to discuss drawing format for reports. Tony Matre’s request for ‘General Information Notes to occur on each floor plan are discussed.

05/31/12 – Draft Report completed.

06/01/12 – Received note from Mark Hesselschwerdt describing final step for process of entering data into Archibus

06/11/12 – GBBN set up conference call meeting with Facilities CAFM Department and CAFM Partner to be held July 3, 2012.

09/06/12 – Bulk drawings received from Pinnacle.

09/24/12 – GBBN completes bulk drawings and forwards to Pinnacle for use in final reports. GBBN requests final inventory drawing spreadsheets to allow completion of inventory drawings.

10/27/12 – Inquiries to Pinnacle Engineering about identification of ACM designation clarification for Archibus entries. All other Inventory Drawing and spreadsheet work completed.

01/26/13 – BRG confirms entry of ACM data into Archibus is completed.

01/28/13 – GBBN conducts spot review and confirms data entry is accurate.

02/26/13 – GBBN inquires with Pinnacle for status of remaining documents to be sent to GBBN.

02/27/13 – Pinnacle reports back that progress has been made and further timetable for receipt will be coming 2/28/13

04/25/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response.

05/29/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule milestone goes without response. Reported to Tony Matre via telephone call.

06/26/13 – Email sent to Pinnacle Engineering inquiring about status of project reports and unmet schedule. ThermalTech modified and sent new scope to GBBN.

08/21/13 – GBBN re-submitted SOSS to Bert Watts.

08/29/13 – Bert reports he has to confirm with the Judge about scope and cost for the latest work. SOSS is signed and we are ready to go but Bert says we have to wait until the Judge gives us her approval to release us to begin.

06/26/14 – Project on hold pending litigation.

08/29/14- Pinnacle Engineering sends invoices for services to GBBN.

10/29/14 – GBBN sends Tony Matre SOSS invoice status with all of the Pinnacle Engineering invoices included.

10/31/14 – GBBN pursuant to Tony Matre’s request sends letter of documentation regarding scope overruns.

02/20/15 – GBBN in communications to finalize the remaining invoices to close out these projects.

by TEI TEI

800 – 12th Floor Renovation Project

10/13/05 – Ben sent out some status notes trying to get Arrow to finalize. We are awaiting further direction for our part of the completed work.

09/29/05 – Arrow said they would address the additional items noted at 800 Broadway.

09/29/05 – Additional items were noticed from at 800 Broadway, “This is an F.Y.I. for you regarding the Judge’s bathroom. We had to call the maintenance folks because we ‘heard’ a funny sound coming inside the little access panel on the outside wall of the bathroom located in our kitchen area. When I opened the panel, water was spurting out from two sides of a pipe. The guys came up and turned off the water, I think. Not sure what comes next. However…they mentioned to the Judge that her bathroom sink faucets are on ‘backwards’. You might want to add these situations to your ‘punch list’?? Thanks.

09/27/05 – Arrow responded to punchlist items. PDF

09/23/05 – GBBN issued a revised punchlist with photographs and a directive to Arrow. PDF

09/22/05 – Meeting / Punchlist. GBBN is typing up additional punchlist items, including past unfinished ones, for Arrow to respond to by 09/26/05. At that time it will be decided whether to let Arrow finish, or proceed with hiring another contractor to finish.

09/13/05 – Meeting / Punchlist was cancelled for today. Diana went over to Job Site and is still displeased with Arrow’s Tile work.

09/09/05 – GBBN has scheduled a meeting at the site for 09/13/05 at 7:45 am

09/06/05 – Debbie Poland informed ThermalTech of Judge Grady’s choice of fixtures, and ThermalTech relayed the information to Arrow.

09/02/05 – ThermalTech contacted Arrow and mentioned Pat Murphy at Select Spec Sales, as a possible vendor to obtain a solid toilet seat.

09/02/05 – GBBN informed ThermalTech that the light fixture for over the Mirror has been discontinued, and to specify another one.

09/02/05 – ThermalTech had GBBN forward the following to Arrow: The cleanout and floor drain should be level with the tile and grouted in closely. It should be almost a perfect fit when done right and have a brass cover plate that should look relatively nice. Arrow should adjust the floor flange on the cleanout and floor drain so they are level, if they still can. They may have to bust out the floor and adjust it and then pour a patch around it.

09/02/05 – Judge Grady has made the following requests: Raise the top of the plumbing clean-out to align with the top of the tile. Install a solid toilet seat cover in leiu of the “open” toilet seat currently installed. Install the second coat hook. Install the over – cabinet light fixture. Replace the cracked tile at the “clean-out”. Patch up the plaster adjacent to the sink and cabinet (at door). Provide a finish on the kick-plate of the millwork. Provide shelf in wall cabinet per sheet A201.

08/29/05 – Architectural and Mechanical Punchlist was completed.

08/08/05 – Project meeting postponed for one week.

08/01/05 – Exhaust Fan is not installed yet but is on order. Plumbing is roughed into the wall through the floor space. Electric is roughed into the walls and ready for inspection, insulation and the gyp board. Progress is slow from the MEP side but is going forward.

06/16/05 – GBBN and Arrow are working to resolve a toilet outlet issue. There is a beam below the slab that is interfering with the toilet outlet. GBBN called today to say that they can use a 4-inch filler space to move a wall and thus give the plumber enough room to route the new sanitary line. TE approved this shift in the plumbing.

06/10/05 – TE received insulation, plumbing fixtures and electrical. TE will send out comments on Monday.

06/07/05 – TE sent out submittals on EF, DWV Piping and Vent, Ball Valves and Hangers.

05/10/05 – GBBN setup preconstruction meeting for project on May 24th.

04/19/05 – It was agreed that since Arrow won every part of the bid that GBBN would write an inclusive contract the entailed all items on this project. TE thinks this is great idea and removes all coordination issues between contractors since we will have only one contractor on this job doing all phases of work.

04/15/05 – Bid Evaluation was released. PDF.

04/06/05 – Add01 was released by GBBN and TE. Add01

03/22/05 – Prebid meeting complete. Reupert was only mechanical at Prebid. Sam filled out permit application for HVAC and sent to Brady. GBBN already has all the books and drawings with our stamp. As usual we cannot stamp plumbing drawings because we don’t hold a Master Plumbers license.

03/14/05 – Client review meeting. TTC coordinate with Brady on plumbing items to state the 3-5/8″ wall is fine for our tank type water closet and countertop lavatory.

03/08/05 – TE asked Facilities for direction on copper vs. no-hub cast iron for the sanitary lines on this project in the even they ever entertain waterless urinal systems.

02/25/05 – TE has begun design the MEP on this project. Layout and specs are being compiled.

02/22/05 – TE provided a price of $100 for the gooseneck faucet and $150 for the additional light above the vanity mirror pursuant to the email from the County today.

02/11/05 – DEW advised TE to write the SOSS for this project and submit it to the County for Approval. TE finished and put in the mail on Monday.

01/31/05 – GBBN assembled a complete cost estimate for this project and turned over to the County. We are awaiting further direction from the County. TE noted that GBBN had $3500 too much because TE had front end cost in the estimate and it isn’t required.

01/11/05 – ThermalTech has conducted all fieldwork and assembled a probably cost estimate based on our understanding of the scope of work. We are awaiting a meeting time or notice to proceed. Estimate includes new fixtures, exhaust fan, ductwork, plumbing and electrical work for anticipated scope.

by TEI TEI

800 – Florida Room Renovation

12/26/06 – TTC has decided to close this project and archive it. This project was never contracted or awarded and the work was never done. We believe it was included in another project.

10/14/05 – TE received word that this project is still a “go”, however we have not received any news about the contract status.

10/03/05 – Contracts are turned into the prosecutor. We are awaiting word from the County about proceeding with this project or canceling this project in lieu of a revised EMU project.

09/22/05 – TTC dropped off the revised contacts to Wayne Trent today and gave one copy to Michael Evans at GBBN. All contracts were hand delivered. The County facility group got to meet Kandy today because we stopped in

08/29/05 – Addendum 01 PDF was issued with the new bid form. Job is set to open on September 1st.

by TEI TEI

800 – EMU Relocation Project

12/26/06 – TTC archived the project.

11/02/05 – Diana called today to say that the new SOSS won’t be used and the old EMU #147 will be used instead as we have $3700 left on this job. TTC and Spencer discussed new layout and Spencer sent over a new floor plan for our use. We are trying to minimize mechanical work in the space.

by TEI TEI

800 Broadway Electrical Device Identification

11/16/06 – NR visited the site to verify the panel names for the Arc Flash Study. All panels are verified and the single line is complete. The Arc Flash Label information is currently being inputted into EDSA and the labels are being generated. This portion of the project should be completed within the next 3-4 weeks.

09/27/06 – The final drawings were sent out to Queen City to be delivered to Bert Watts so that he could distribute them to the appropriate people. Areas on the 4th and 5th floor will be updated when the Adult Probation Renovation project is completed and ThermalTech receives the as-built drawings. The panel schedule binders were delivered at an earlier date. All other work on this project is complete.

06/27/06 – The close-out meeting was held on location. There were no open issues. Ginter Electric provided (2) CD’s with the panel schedules on them. I will have the drawings together within the next couple of weeks. TE will have (3) full sized sets delivered to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Darnell Edwards). TE will make (3) binders of the panel schedules and deliver them to Hamilton County Facilities (Attn: Bert Watts, Jim Eby, Darnell Edwards) .

06/21/06 – Spoke with Darnell Edwards about the EMU area. He informed me that the devices were marked and that he is satisfied with the work performed. Close-out meeting has been schedule for June 27 at 9:00am.

11/15/05 – Contracts have been submitted.

by TEI TEI

800 – JV ISSUE DESK REMODEL

03/20/07 – Check CRA website for closeout documents. Project is ready for archiving.

02/19/07 – Coordinate with CRA for final details. Project is nearly complete.

01/22/07 – Permit drawings have been submitted to the IBI

01/02/07 – TTC and Spencer discussed the project today and TA and TTC will kickoff the project internally on 01/03/07 and push it towards compeltion.

11/10/06 – They installed the desk. Flooring will be installed in the next several weeks.

09/08/06 – Preconstruction meeting held. Things went as planned.

04/05/06 – Review meeting is scheduled for this Friday. Cost estimates have been given to the client. No HVAC or Plumbing work on this job.

03/08/06- Mike Stiles met with the architect on this project for kickoff.

by TEI TEI

800-York Control Panel Replacement

04/26/07 – This project is complete and ready to be archived per the County Partnering Meeting.

01/05/07 – TTC has given the County everything they need to get the PO for this work and it will be done in house on this project. York has provided all the paperwork for sole source and quotes.

10/25/06 – Prebid Mtg today. DK showed up for prebid. TTC built the quotewire package and sent it to purchasing for bidding out just the purchase. TE has to send out an Addendum for bidding changes.

06/27/06 – Tony stated for us to get the electrical requirements together for the trades so they can do the work. TTC said he would get with TA and get this done.

05/05/06 – Mike D and TTC discussed recent information provided by York. Additional scope items and more electrical work that would have to subbed out on the job. TTC talked to service tech’s about doing this work and is trying to formulate a complete scope for the project.

04/26/06 – TTC reviewed the project scope with Tony, Mike Denton, Darnell Edwards, Bill Scholl and Tony Martini. Tony need further proof that this project should not be outsourced because o requirements on County Efficiency Committee.

by TEI TEI

800 – Security Camera Upgrade

Journal:

11/29/07 – Bert reported today that final pay application has been submitted for. This project could officially be closed out at this point. Bert has confirmed that the County has received the O&M’s and As-Builts sent by TTC at TEI.

11/16/07 – Lori contact TEI today about past due notices. TTC collected information and sent to Nor-Com saying that TEI would do as-builts and write deduct change order of $2000 for the job and Nor-Com has to provide a full warranty letter to HamCo. TEI finished the as-built and made a CD of the dwg. TEI finished the change order for a $2000 deduct. TEI sent both to Nor-Com to return to Bert Watts with the final warranty letter. TEI finished the as-builts and change and mailed to Nor-Com and the County. We are finished with this project.

10/18/07 – TTC opened and sign payapps 3 & 4 and sent final closeout paperwork to Bert for processing. This packet arrived sometime after 09/07 with the wrong zip code. All the certified payroll, closeout documents and payapps appear to be correct. TTC emailed everyone and said the job is ready to close.

06/07/07 – Lori from Nor-Com checked on status of payapps. I emailed back #1 show paid, #2 i need 3 originals and #3 i do not have a copy of.

05/31/07 – Closeout documents have not been sent. They were waiting to use the allowance money. MD directed Nor-com to close the job in an email and TTC forwarded that email to Mike Rolf.

05/30/07 – TEI does not believe they have received close out documents yet for this project. TTC contacted Mike Rolf.

03/30/07 – Punchlist is complete. There are no items from TEI or the owner at this point.

03/23/07 – TTC was on site today and took down punchlist items for Nor-Com. Currently there are no punchlist items for Nor-Com to correct.

03/15/07 – Nor-Com reported that cable work has begun in the space and that unit operation training is complete. 18 people attended including Phil Lambing from TE. Nor-Com will provide a more concrete completion date early next week.

03/14/07 – Nor-com was scheduled to begin work today and Vicom held their master training session for County employees this afternoon.

02/19/07 – Nor-Com is sending a new CD for submittal review on this project. TTC could not find the original.

02/08/07 – TE distributed Const Mtg notes for the pre construction and Const Mtg 01. TTC did not make the meeting in person but attended via a conference call.

02/01/07 – TTC sent MWH dwgs of the space for his use.

12/28/06 – TTC discussed kickoff meeting for this project and sent out another date because all previous dates have conflicts. We are now trying to have a kickoff meeting on Jan 4 at 9am at the 800 BWAY building. Norcomm has already sent in RFI’s on this project and proposed a solution to a DVR problem by downsized the 900 GB harddisk to allow for multiplexing of the signal to the DVR. This will have to be discussed at the kickoff meeting.

12/26/06 – TTC created project journal and filled in back data from previous meetings.

by TEI TEI

800 – ADULT PROBATION RENOVATION PROJECT

Journal:

11/16/07 – TEI received a hard copy of the approved QCE change order in the mail today. TTC filed it for archiving with the project folder. Mike Cavanaugh called to inquire about payment of this invoice and TTC referred him to Bert Watts because all invoices should be paid on this job by this point. TTC is ready to archive this project.

11/06/07 – CRA forwarded Change Order #2 to Bert Watts.

10/30/07 – CRA issued Change Order #2 to delete the allowance balance and forwarded to Kramer & Feldman.

10/26/07 – CRA received final closeout documents.

10/18/07 – MD sent back final change order for signatures on 3 originals. TEI printed and sent 3 new originals to QCE for signing.

10/04/07 – TTC and LKV revised CO4 and sent it to QCE for signing. The new amounts were gleaned from the previous emails and signed change orders.

10/01/07 – TTC received, reviewed and signed payapps #5 and #6 from QCE for this project. I read through all the email correspondences and believe these final two payapps to be correct. Assuming that the change orders that the owner had in hand were signed and correct, the amounts of these pay applications should correspond to requested amounts by the Owner. TEI will mail them to Bert Watts for processing today.

09/28/07 – Awaiting final closeout documents. CRA needs As-built drawings.

09/04/07 – Items mentioned in previous entry were sent to QCE this morning.

08/31/07 – Pay application #05 and #06 need to be updated to show CO#04 instead of CO #05 in the work description (there were only 4 CO’s); the dollar amounts shown are accurate. TEI will correct the dollar amounts shown on CO#04 and send 3 originals to QCE for their signature. TEI will also send 3 original copies of allowance clarification #01 for the entire allowance ($2500) to QCE for their signature. When QCE has received and signed all 6 documents they will send them along with the updated pay application #05 and #06 to TEI. All of the above mention documents will be reviewed, sorted, and sent to HCFD. The documents will be sent to QCE on Tuesday, September 4.

08/23/07 – TEI printed 3 more originals for Change Order #5, signed them, and sent them to QCE with instruction to sign them and forward them to HCFD. TEI also sent the O&M manuals along with the final pay application and closeout documents to HCFD Attn: Bert Watts.

08/17/07 – TEI is awaiting the O&M manuals from QCE in order to close this project out. NR contacted Greg at QCE on August 16, 2007 in order to locate these binders. Greg said he sent them to TEI already but we have no record of receiving them. NR will contact QCE today and have them send another copy.

07/31/07 – TTC approved Change Order and PayApp # 5 for this project but asked for backup data for the change order and requested

07/30/07 – QCE is not responsible for completing the ceiling patch work. The remaining ceiling patch work is going to be completed by the Hamilton County trades department. QCE has completed their portion of work. TEI is awaiting the resubmitted final pay application and close out documents in order to close this project out. TTC emailed the County and QCE for closeout documents on this project and approved their last payapp and told CRA the same info.

07/09/07 – TE still awaiting modified change order from QCE.

06/29/07 – QCE has been informed that they will not be paid for the ceiling and wall patching. TE is awaiting a revised change order and final paperwork from QCE.

06/15/07 – Still awaiting response from QCE.

06/15/07 – TE placed call to Queen City Electric about inaccurate Change Order #04. Awaiting response from QCE.

06/08/07 – Change Order #04 was sent to Hamilton County Facilities Department but had inaccurate information. The change order did not reflect the three previous dollar amounts. The paperwork was given to TA of TE in order to straighten the numbers out.

06/01/07 – TE is still awaiting response from contractors about the punch list. New contract dollar amounts between the contractor and Hamilton County Facilities Department are being clarified. This should be resolved in the next couple of weeks.

05/25/07 – TE issued punch list. Awaiting response from contractors stating that all items have been completed.

05/21/07 – TE completed punch list. Issues with lamps have been corrected. Punch list will be issued early this week.

04/27/07 – Fixtures are 99% (one recessed can yet to be installed). TE has visited site on 2 occasions this week. Project is expected to be complete and TE finish their punch list by the end of next week, May 4 of 2007.

04/20/07 – Project is moving along as expected.

04/13/07 – Light fixtures are currently being installed. Project should be complete by the end of the week of April 20 of 2007.

04/09/07 – Phil stopped down at the site on Thursday and resolved some clarification issues on the type of lights the supplier purchased for this job. TE instructed them to send them back and get the correct cuts for remodel 5″ cans to fit into the drywall soffit.

04/05/07 – QCE has begun installing the fixtures into the space and then we can close this job out.

03/30/07 – Fixtures are scheduled to be installed next Wednesday, April 4 of 2007.

03/23/07 – Project progressing as expected.

03/20/07 – Lights are scheduled to arrive next week and be installed next week by Queen City.

02/19/07 – TEI is waiting for Mindy to get back for funding approval but lighting changes are happening now. Check CRA website for details.

01/10/07 – Const Mtg was held but TEI wasn’t in attendance. Project updates are available on the CRA website.

12/26/06 – TTC moved project over to Mark Donnelly and will get an update from Phil and Spencer this week.

02/06/06 – Spencer submitted a revised floor plan with a new sink location in the back corner of the building. TTC and SBO will pipe up the sink and drain while PEL finishes the electric work. We are shooting to have this job on the street by the end of February 20th for this job.

08/04/05 – Had design meeting today. Discussed schedule and typical workstation layouts, phasing of the project and construction times. C+R will issue official meeting notes for the job. They are talking about moving in around November so design documents have to be collected pretty quickly.

06/24/05 – TTC and Spencer visited the site today to look at the lighting and ceiling heights. TTC brought pictures back for Phil.

06/16/05 – Cost estimates were given to client and we are awaiting direction from the County and C+R before proceeding. TE hasn’t heard anything since 03/21 on moving forward with this project. NO HVAC or Plumbing were included in the cost estimate to the client.

03/01/05 – SOSS Written. TA to approve and send to DES.

by TEI TEI

800 – VACUUM PUMP UPGRADE

07/09/08 – TTC stopped by the site today and review the project progress with Process. There were several issues with the installation of the equipment that we resolved. Size, location, pipe runs, housekeeping pad, inlets.

12/04/07 – We received a schedule of values today that TTC marked up and sent back to Process. Insulation submittals were received today in the mail. LKV is logging them in for this project and will give them to TTC to review.

11/29/07 – Joe Doxsey and Terry Cannon discussed project schedule and submittals today by phone. Process is going to send their background checks to Willie at the County and are currently working on the submittals. They have not worked on submittals or project schedule yet according to Joe but will get them done now and into the system as discussed with TEI today.

11/16/07 – We held the pre construction meeting with Process and discussed changed scope and timetable for the project. Bert provided an email with the PO attached yesterday. TTC asked for the submittals on the vacuum pumps for the job and saw Rick Huff today and asked him if they were ready. He had no idea about them being asked for so Process must not have told him yet. Laurie finished typing the meeting notes and TTC has to review them and send them out. Process is going to provide a project schedule and SOV for our approval and ThermalTech has printed out and is sending 9 sets of stamped permit drawings (6 for city, 3 for IBI) for Process use on this job. TEI also emailed Process a copy of the spec and the dwgs for their personal use and printing.

08/02/07 – TTC sent out Addendum 02 for the project today PDF. Addendum 01 extended the bid opening to August 7 by Purchasing.

07/31/07 – Bernie and Roger replied with Addendum 01 items for this project. A&M stated that their cost would be about $4500 for this work even though we are adding about 8 points and picking up alarms for the system. Bernie added that the 3″ line sizes were all he needed from the dwgs.

07/30/07 – Purchasing issued Addendum 01 for this project extending the Bid Opening and Due Date to August 7, 2007. TTC will issue Addendum 2 with pipe sizes, control notes and spec revisions.

07/16/07 – TTC sent the same data to Queen City to hold for bidders this morning. TTC has to contact Gina about the advertised price for this project since he changed it after talking to Tony Matre.

07/13/07 – TTC sent the project out for bidding with a bid opening date of July 31, 2007. Rod, Jim and Michael helped complete the project and send it to Gina by 4pm.

06/26/07 – TTC and Rod are going to finish the mechanical design on this project. Michael Huntwork is going to finish the electrical work on the project and TTC has to get project number from Purchasing for this project.

05/30/07 – TTC selected pumps and is finalizing the dwgs and specs for bidding. TTC asked to add the dampers to the back of the combustion air intake while A&M was doing the control work on this job.

05/14/07 – TTC and NR worked on CAD layout and background drawings. TTC will try to have something to review by this Friday. TTC talked to Tony today about using 2 full size vacuum pumps instead of 3 separate pumps. If this works, Tony did not seem to have a problem with the new design.

05/03/07 – TTC visited the site for field work and equipment locations. Meeting notes were distributed to Owner.

04/13/07 – We had second Design Mtg. 3 vendors were selected. We filled out the BOD and DID for the project.

03/15/07 – Tony said the SOSS is approved but TE has not seen it yet.

by TEI TEI

800 – BOILER REPLACEMENT PROJECT

01/22/09 – After making it most of the winter with little problem, the boilers started performing badly last Friday. ThermalTech was on site with the Duke employees to monitor the situation. TTC asked Stoermer and Duke to provide solutions, TTC believes our two options are a new regulator in the boiler room for 14″ or modified regulators at the boilers to handle the boiler pressure better.

03/09/08 – Duke installed a new regulator and it worked good until last Wednesday when there were three low pressure alarms and flame outs on Boilers 2 & 3. TTC sent an email to Duke and SA explaining the latest problems and asked to bring in another expert to look over the entire system.

02/29/08 – Duke installed the 461-S regulator this week and if this works, they will install a larger vent for the regulator to use. We will monitor the boilers over the next few weeks and see if this is the answer.

02/06/08 – Duke Energy replaced the regulator yesterday with the new CL-38-2 but did not use the pilot line. The regulator seems to work okay and settles down in about 2 seconds of a boiler start. TTC sent an email to everyone outlining the results and asked SA to program the boilers one more time with the new regulator in place. The residual static is now up at 47″ with a running static of 27.6″wc. TTC witnessed a dip in pressure down to 17″wc with the new regulator but it recovered quickly and evened out as the boilers started.

02/01/08 – We had a conference call with Duke Energy today. They are willing to try a new gas regulator and meter at the building. We are going to try a CL-38-2 regulator with the new meter next Tuesday at 9am when Duke comes out to install the new meter set. Darnell will coordinate the visit and TTC will stop by to check on progress around noon. They think this meter with the external sensing line and 1.25″ orifice will do the trick.

01/24/08 – Boiler #3 went out on a flame failure last night at 10:30pm. John Gilbert came in to reset the boiler and it has been working fine since then. TTC had emails with Duke & Stoermer to try to get these final issues resolved.

01/23/08 – On 01/18 SA and Duke met on site to work and reset the boilers. What we discovered was that Duke was not comfortable with having the pressure under 1psig because we were testing the boilers at 24 psig and 20 psig. At these pressure the boilers worked pretty well but there are still concerns with Duke regulator. Because of this TTC talked AKM into paying PHB to come on on Sa 01/19 to install a 1″ weld-o-let and valve so that Duke could move the sensing line downstream about 20ft. On 01/21 TTC met Duke back on site and they piped up the regulator sensing line. The regulator is still surging. The boilers are however running great. TTC check the pressure at the inlet to the boilers and it is bouncing around between 19 psig and 29 psig every second. The burnertip pressure however is stead between 7.8 & 8.0 inches. I believe the biggest change in the boilers is that TTC talked SA into changing the low pressure setpoint down to 13″wc and not 19″wc so the boilers are not going out on low gas pressure as much any more. The Sellers regulators seems to be able to handle the incoming gas variations very well and the boilers are firing very smoothly.

01/15/08 – Darnell Edwards called yesterday to say they were having more trouble with the boilers. I met Duke Energy on site today because they responded immediately to a request to test the gas pressure at 15″ wc. It turns out that the boilers can’t be run at the pressure because the low limit on the boilers are set to 19″wc for some reason. Duke set the regulator down to 24″wc incoming for now and the boilers seem to run fine during the day. I am waiting to see how they operate overnight and tomorrow morning. I have talked to Tom Schmidt and Joachim at Duke about this situation too. We are currently trying to schedule Duke and SA back out at the building at the same time to adjust the incoming gas pressure to 20″ wc which should put us right in the middle of the operating range for this boiler. I am also trying to get some chart recorders to log the gas pressure in the main and at the boiler burnertips for testing purposes. When we watched the pressure after the regulators today they were steady at 7.5″wc on BLR-1 and 7.9″wc on BLR-2. BLR-3 is having some startup issues and according to Tony Martini there is sooting on the fire tip inside the boiler. When the boiler cools down, he will clean it off and then it should be back in server as well. I am trying to coordinate a meeting for Friday morning at the 800 Broadway building between TEI, Duke and Stoermer. I am also trying to get chart recorders for history recordings of gas pressures.

01/07/08 – The County called last week with more boiler problems. TTC investigated and found that on November 9, 2007 that the boiler were setup for 37″ of gas pressure on Jan 3, 2008 during our first cold snap the boilers were not running properly and going out on alarm settings. The owner could not keep a boiler running. TTC visited the building on 01/03, 01/04, 01/05 & 01/07 to check on the boilers and the pressures in the main gas line. According to all the readings that TTC took while onsite there is no gas pressure issue with the natural gas piping or Duke regulator. During all test the natural gas system delivered 30″ of pressure to the boilers all the time. At startup or when both boilers were starting at the same time a dip down to 24″ could be seen for a fraction of a second but nothing that could be thought to effect the way in which the boilers are running. TTC, John Gilbert & Will check the boilers at Boiler #1 and at the main building entrance. There is no appreciable loss in the piping system in gas pressure. The new Duke meter seems to be controlling pressure well. Today the temperature was back up to 60 degrees and TTC tested the natural gas pressure with no boilers running and it was found to be at 30″ with both of the domestic water heaters running. Later after they both shutdown the pressure slowly rose to 37″ just like the Duke representative Mike Hoffer said it would. The regulator used back pressure from the high pressure main to ensure positive shutdown of the valve. This means the residual pressure in the gas main is 37″ when no equipment is running and 30″ when anything is on. This correlates exactly to what to Ron Eaton told us by phone last week. The next step is for Sellers to provide a new regulator on one boiler sized for 30″ of static pressure when the boilers are running. This regulator should be able to handle pressures up to 40″ of gas and down to 28″ nominal and even down to 20″ under heavy load conditions or even better if it can go down to 14″, that is what I will make them provide. In my opinion, this is absolutely a boiler regulator issue and not a Duke Energy gas issue. We can hopefully get a new regulator in this week and have ONE of the boiler retrofitted before the next cold snap but your also stand a great chance of the boilers working well right now since they have been adjusted down from 37″ of pressure to 30″ (I think that is just a misunderstanding on Stoermer’s part about what they have at your building).

11/09/07 – Tom Schmidt called today to say that the boilers could not be adjusted so that all 3 boilers run at once. TTC sent an email to the County and Tom stating that all 3 boilers can’t run at the same time. The design is to have one boiler running at a time, at 0 degrees a second boiler can come out to help out but it doesn’t have too. The third boiler does not come on until -20F which really means never because the 6″ gas main isn’t sized for (3) 125 HP boilers to operate simultaneously. TEI had Tom Schmidt come in to discuss the problems and talked to Ron Eaton after he got to the building today about inlet pressures. On Nov 9 the pressure was 37″ and the boiler were adjusted for that pressure. Today the pressure is 29″ and this changes the inlet pressure to the boiler after the Siemens gas train from 8.5″ to 7″ which causes the boiler to run way to lean. Gary and Todd are helping me figure out how to solve these problems for the County. TEI has a conference call with the boiler mfgr at 2pm to work on this some more.

05/19/07 – Duke Energy replaced the gas regulator to the building with the new model specially designed for our load with a tighter tolerance to the outlet pressure.

05/02/06 – Tony called today about flame failure on the boilers at the end of the last season. TTC recapped what has transpired since the 3-22-06 call. TTC has talked to both Stoermer and Cinergy about this problem and it appears to be rooted in the fluctuations of gas pressure not the absence of gas pressure. The building has at least 1 psig of gas pressure which is twice as much as the boilers require but when boilers are running with the water heaters and various units shutdown, there are flame failures in the system. Apparently, the fluctuation of gas pressure on the high side is causing gas delivery problems to the boilers. TEI may recommend having the water heaters moved ahead of the boilers or changing the control sequence so that 3 boilers are never online at any one time. TEI will contact Stoermer, Cinergy and Automation to make sure this problem is resolved going into next heating season. We might need a meeting between TE, Cinergy, Stoermer and A&M to resolve this issue.

03/22/06 – County called about Boilers 1 & 3 going out on flame failure last Thursday and Boiler 2 going out on Saturday.

10/28/05 – County called about erratic boiler controls for gas usage. SA worked on setting the boilers up again with Cinergy regulators.

06/14/05 – TE received and sent out all final paperwork to complete this project for County and PHB.

06/09/05 – TE and PHB discussed final paperwork requirements on this project. TE received a final invoice in January 2005. The final paperwork was received in January also. Now, TE has all the final paperwork in house but can’t find the final application # 5, so PHB is sending over a new original. TE will submit these final documents to County as soon as they are received, copies and signed.

05/03/05 – A&M straightened out the program and released some unpronounced overrides. The provided a little more training for the operators and building manager. The plant appears to be working properly.

05/02/05 – 800 reported some anomalous behavior from the boiler plant. TTC called A&M for dial in assistance and TTC will stop down at building on Tuesday morning to investigate.

04/27/05 – TTC and Tony Martini witnessed the laptop working remotely today. Dennis set it up for home use and left it for Tony to take home for use. TTC observed the entire system while the Free Cooling system was offline because of a condenser water pump rebuild. I found many things in the “override” position. This certainly happens upon occasion but overrides should be release to allow the automatic controls to handle the building most of the time. TTC pointed out that with the boilers overridden off at the DDC through their enable point that this does not disable the feedwater pumps so we actually had 0 boilers running but 3 feedwater pumps running just using up energy. TTC stated that a better way to control the boiler plant “off” would be to lower its enable setpoint below the current OA temperature. The same goes for the Chiller plant which can also be run at lower temperatures when the FREE cooling is disabled to keep temperatures in the building from overheating the space as they were beginning to do today.

04/26/05 – TTC could not get the laptop working from home. TTC delivered to Dennis Brown. A&M called to report the laptop is working and will be dropped off to 800 on Wed morning with training. TTC will stop by to supervise this training. Does the County want Tony Martini and Glenn Wilhoit to attend this laptop training and can they be there at 8:15 a.m.?

04/22/05 – TTC to work on laptop for DDC access. PHB is complete on job. Ron Eaton with Stoermer Anderson made some boiler adjustments today. Cinergy increased building pressure to 2 psig at the meter and in the building. SA said boilers are working fine and building staff is happy with performance at this time. SA also said that PHB had a regeneration line from the water softener to a floor drain that still has to be piped. Everything else appears to be complete.

04/18/05 – TE contacted A&M about 800 boiler items that need to be worked on such as Laptop control from remote location, outdoor air sensor is not calibrated and 2nd floor ventilation CFM issues. TE dispatched SBO to assist in troubleshooting the airflows on the second floor. Sam recommended a change in airflow deflection and balancing and further suggested relocating one of the airflow devices away from the center. Watch changes made and notify TE if you would like to proceed further. A&M will schedule onsite time with John Noyes to work on OA sensor and laptop programming.

03/24/05 – Tom Schmidt reported that all devices are up and running with the Boiler system. Ron reported that a setting was changed on the softener system and that is now running also. This should make all items complete for the 800 Broadway Project.

03/11/05 – PHB delivered the O&M’s to ThermalTech for this project. TE will review and turn into the County for record keeping.

03/07/05 – TTC contacted all contractors to resolve issues at the 800 Broadway Building. (1) Trap on boiler # 1, (2) Sample Cooler on Boilers, (3) Low Fire control on boiler, (4) Draft controls on boilers.

03/01/05 – Tom Schmidt reported that 2 of the 3 draft controls are upgraded and working. The third one has a bad magnahelic. It will take one week to get a replacement mag from the vendor. SA will return to install the final automatic draft control components at that time. PHB will have a pipe fitter to pipe in the replacement sample cooler within a week according to PHB.

02/25/05 – PHB will have a pipe fitter down to install the sample cooler and remove the outlet drain valve within a week. TTC is trying to find out the status of the draft controls upgrade.

02/15/05 – O’Shea patches the concrete yesterday. SA is supposed to be installing the draft controls today.

02/11/05 – Stoermer dropped off the sample cooler yesterday. TTC will contact Peck about installing it. TTC will have Automation setup training next week and upgrade Tony Martini’s Laptop. TTC contacted SA about a $300 service ticket for the water softener as we believe this is only

02/10/05 – Stoermer reported today that the boiler draft controls will be installed on 2/14 or 2/15 with Ron Eaton. Tom Schmidt will personally be delivering a new sample cooler today to the building and he also stated that this type of failure is normally associated with the cooling water to the cooler not being turned on quick enough and an over-pressurization occurs in the shell vessel. This may have been what happened here? Who knows. To be safe, I would advise opening the drain first, then the cooling valve and the test water valve last? I am sure Tony Martini already does it this way anyways, but just in case, I thought I should report what was given to me.

02/02/05 – 800 reported today that the new softener cannot keep up with the water load into the CR tank. TE is investigating this situation and will advise soon.

02/01/05 – TE request updates from contractors for these issues: Water Sampling Station, Draft Controls, Steam Leak at Flange (up high & extra charge) and Concrete Repair.

01/26/05 – TE faxed request for final documents to the attention of Amy Reeves of Peck Hannaford & Briggs. TE coordinated with PHB on fixing concrete flooring in boiler room. TE to discuss with County painters about re-epoxying the floor to match existing red floor previously installed.

01/19/05 – TE faxed back seal pump replacement quotes to Tony for Chiller Room.

01/18/05 – TE talked to Stoermer Anderson about wiring the drafts controls and we still haven’t received new wiring schematics. Tom Schmidt convinced them that just wiring schematics isn’t what we need, so they are sending new internals to the site for installation with new operator manuals that will be up to date and correct so that in the future the wiring schematic will match what is inside the draft control boxes. This work, however, won’t happen until next week at the soonest.

01/03/05 – TE trying to finalized punchlist and get control system backed up on CD format for County. PHB to begin final punchlist work and SA to retrofit draft controls this week according to Tom Schmidt.

12/13/04 – 800 called today because unit heater was not running in electric room and the space was cold. TE advised that the heater should come on anytime the sensor on the interior wall reached 45 degrees. Also, there is no graphic for this heater or control over this heater from Metasys. TE will stop down to check on situation today and help out.

12/10/04 – TE confirmed that SA should be on site on 12/14 or 12/15 to fix the water softener and the draft controls for the boilers. TTC told the building staff that the OA setpoint for the boilers can be used to adjust the start temperature of the plant at anytime. 800 reset the temperature up to 55 degree F.

12/02/04 – TTC stopped at building today and observed that boiler 2 is not running well. He sent emails to contractors for repairs. TTC noted punchlist items that need to be completed and contacted PHB about schedule to finish.

12/01/04 – Automation made final control changes at the site today to fix runtime accumulation modules, labeling the DX-9100 modules and verify the operator control points for lead and lag boiler start. TE forwarded the email received in response to that visit. A&M will update the sequences and get as-builts drawings to TE for inclusion in the final O&M’s for the County soon.

11/26/04 – TTC stopped at the building on Wed (11-24) and Fri (11-26) to verify and watch operation. The boiler plant is running in complete automatic operation at this time and ran appropriately over the Thanksgiving holiday. On Friday, the building was warm and there were no critical alarms on the boiler plant or any manual resets required at either the boiler of the condensate receiver. Todd Zurweller provided the final piping punchlist that will be converted to as-builts drawings that will be turned over to GBBN for inclusion into the CAFM sets. TE will follow up with Peck and Stoermer on the completion of punchlist items within two weeks and TTC will follow up on some control issues and boiler ops issues that he would like to see defined a little more clearly. SA completed the replacement of the solenoids valves, additional strainers in the feedwater system, HOA switch on Boiler #2, and new overloads on the all 3 boilers to prevent nuisance inrush tripping. SA still has to repair the water softener control plate assembly and replace the internals to the draft controllers for all 3 boilers. TE to expedite the remaining items. Remaining items to complete on this project (to TE’s understanding) are: CR water temperature at 150 degrees F (design calls for 190 degree F (SA shall adjust control valve to preheat the feedwater to meet spec), water softener repair, draft control replacement, sump lid replacement, seal replacement on vacuum pumps, seal replacement on cooling plant pumps, concrete repair in boiler room, A&M has to fix the runtime hours display, A&M has to label the DX’s inside the panels for points, A&M has to provide updated as-builts for controls (TE reviewed preliminary submissions with Roger Corbett on 11-24), Peck must install breeching cover must be installed, Peck has to replace or fix the trap on B-3 as it blows by when in normal operation. Glen Wilhoit found a bypass valve open on the condensate receiver, which when closed, fix the steam blow by problem into the floor drain behind the CR unit and stabilized the water level inside the CR immensely, TZ commented that the trap on the CR unit overflow meets the 12″ dwg requirement but TE prefers this trap to be as deep as possible as this is what provides the non-mechanical seal on the CR tank to keep any steam from escaping (we will check back at end of project to see about allowance monies left to increase depth). Peck shall have insulator finish up all insulation on project (lots of insulation work left to do). TE instructed SA to disable the high water level alarm on the CR tank because of nuisance tripping (this alarm is for notification only and requires no user input to fix anything so we just turned it off – the staff reported that the old tank would have high water conditions too but never had an alarm so we will just match the old operation, the tank itself is sized for 4 boilers condensate return so it is of ample size for this plant per Todd Zurweller. The water TDS from the sample coolers has been reading like this this week: at the CR 600-900 micro siemens (this makes sense, city water is about 300-600 normally), the boiler samples have shown 3.6 milli siemens or 3600 micro siemens which is also right in line with what we would want (i think 2500 – 4000 is okay, anything in the 6000 or higher would tend to lead to carryover) and lastly, TTC adjusted the gate valves on the outlet of the boilers to about one hands width of stem showing in an attempt to calm down the surging of the boiler sight glasses, this seems to let the water level stay more consistent and allows the make-up water to keep up with the boil off rate at all conditions, the closure of these valves may be changed when the boiler plant is under heavy load if needed.

11/23/04 – TTC watched boilers all day last Friday. The water controls are such that the boiler plant is too finicky to leave in total automatic operation without expecting nuisance alarms from either low water at the boilers, high and low water at the condensate receiver or overfilling of the boiler vessels through the solenoid valves. SA reports that all components replacements are scheduled for Tuesday except for the draft controls which will be done ASAP. TZ is sick so the TTC & TZ field trip has be rescheduled until Wed at 2:30 p.m. If Glen Wilhoit has any questions please have him call TTC at the office. TTC discovered that the solenoids are letting water bleed past even with a pump on all the time and not the check valves as the water level stayed the same all afternoon with the feedwater valves blocked. In addition the traps on the boilers only drain the boiler vessel when there is a positive pressure inside the vessel, once in a vacuum they are not draining enough water to keep the boiler from overfilling on this day.

11/22/04 – TTC talked to Glen today and we discussed looking at these issues: slope of main steam header towards traps, conduit cover on Boiler # 03, Vent on receiver tank. TE will investigate these issues on Wednesday at the TZ walk through.

11/18/04 – Boiler #1 is overfilling with water. Last time this was due to dirt in the feedwater line blocking the solenoid valve. When the new valves and strainer are installed by SA, this problem should be eliminated. 800 could try to run Boiler #3 instead of #1 to alleviate this problem.

11/16/04 – The boilers are carrying over on startup and shutting down on low water forcing the electric overloads to cut out. SA has to replace the overloads on each boiler to correct this problem, replace the HOA on boiler #2, upgrade the draft controls and install cycle timers into the boilers. This work should be done in one week. All 3 boilers are available for use. At morning startup, Tony Martini may want to turn 2 boilers on and back down the gate valves to 30% of so to allow the plant to come up to working temperature to smooth out this hiccup. Very soon, the boilers can be left on full time. TE is worried about the surging in the CR unit and boiler. The water levels are bouncing around more than we would expect. TE to investigate more.

11/12/04 – All contractors and TE were on site to finish up the project. TE hired Geiler directly to replace the Condensate #4 pump seals because Peck could not meet the schedule. Upon disassembly we discovered that both seals needed to be replaced along with one motor bearing and one motor winding. TE authorized this work to be completed so that the plant can be ran in full automatic on Friday 11-12, Peck service was on site to replaced six sets of seals on condensate pumps CR-1, CR-2 and CR-3, Peck was only able to replace one seal on CR-2 (this work was being done at T&M, TE recommends that we re-evaluate having Peck do all the pumps as we don’t believe there will be enough money left in the allowance to pay 2 days of T&M for each of the 5 remaining pumps, we bought all the seals already which cost $1200, Peck and Cinergy upgraded the gas supply line to 1psig pressure, tested it, installed new regulators and put the system into service. 800 now has a new regulator for the water heaters and for AH-43 on the roof. After purging the system, all the equipment was back in service and burning properly by the end of the day on Friday. We permanently disconnected the garage AH unit that was abandoned, the County CANNOT use this unit unless another regulator is put in the gas supply line (this shouldn’t be an issue since this unit hasn’t been used in 10 years – the gas line has been capped. Stoermer Anderson (SA) finished wiring the draft controls and adjusted all the burners to spec. The draft controls need some different components in the upcoming week as they only work in the wide open position right now. The water softener needs to still be replaced since the Crown Engineering unit was damaged in shipping. The sparge tube and control valve were installed by Peck and adjusted by SA. The unit heater was installed by Peck, wired by Glenwood, and wired by Automation for controls. The unit heater and the exhaust fan worked per spec. The punchlist was commissioned by TTC and will be sent in a separate email where TTC will document remaining items including training which should be given Mon, Tue or Wed of the upcoming week.

11/10/04 – Glenwood will finish up the electrical work on Thursday. Final inspection scheduled for Monday. Please make the accommodations to allow the inspector into the building. All contractors are scheduled for work on Thursday as well. TTC will call automation when enough of the boiler is complete to allow control system to take over control of the boiler plant. Final commissioning and punchlist will happen in afternoon on Thursday.

11/08/04 – All gas pressure work is scheduled for Thursday 11-11-04. The building will be shutdown for Veteran’s Day. Peck, Stoermer, Automation and Cinergy will all be onsite connecting and verifying all final workings of the boiler plant. Terry Cannon will also be onsite to assist in the last minute details. Water softener and unit heater are on site ready for installation (softener has ctrl panel issue). Stoermer has sparge tube in town and hoping control valve will arrive on Tuesday. Stoermer needs to finish wiring the remaining local controls on the boilers. Today’s onsite work proved that one boiler and both water heaters cannot run at the same time without the boiler going out on low gas pressure. PHB assured us that the boiler room door will be installed on 11-09-04. The vacuum pump seals were ordered for overnight shipment and should arrive on Tuesday to Peck. Geiler will be in at 7:30 a.m. on Tuesday to install the seals on the Condensate Pump-4 in the chiller room (ttc coordinated with Geiler and Martini for this work). Outdoor louver needs to be installed. Final insulation needs to be installed. Final control programming needs to be implemented. Training for Tony Martini needs to be provided by Automation (ensure the whole system is demonstrated, optimized control of cooling towers, free cooling, boiler sequences, etc).

11/03/04 – TE confirmed today that the Seller’s boilers can handle and are built for 10 psig inlet gas pressure, at worse case a spring needs to be changed in the regulator and SA will provide that service at no charge. TE instructed PHB to buy two regulators at this point: (1) for the water heaters sized for 2 x 199,000 btuh, 1 inch, 1 psig to 9″ w.c. and (1) for the AH-43, 1 x 160,000 btuh (5″ w.c. nominal, 14″ w.c. max), 1 psig to 9″ w.c. Emergency pressure at inlet to be protected up to 10 psig. As soon as we get the new regulators in we will schedule the work and the shutdown and as TE understands it at this time, Tony Matre has instructed PHB to do this work after normal hours or on a Saturday.

11/02/04 – Remaining project items: insulation, water softener, sample coolers, door installation, damper controls, large louver for combustion air, unit heater, sparge tube, boiler #1 falsely enabled, new regulators for water heater, coordinate pressure test, adjust fuel air mixture, sump lid, chain link fence, seal barometric damper, seals on vacuum pumps, reinstall exit sign. Be complete by November 12 including controls.

11/01/04 – PHB coordinated with TE today on boiler requirements for new venting. Sizing requirements are 17,258,000 btuh connected load (5,620,000 each boiler, 199,000 each water heater, 1 psig inlet pressure (with an emergency inlet pressure of 10 psig), 15 inch w.c. outlet, if a 2″ regulator can meet our flow demands, we won’t need a separate relief valve for the regulator (provided it uses rubber seats), Dead End Lock Up Model type, Main burner and pilot applications, static pressures will be allowed to rise slightly above setpoint, an additional regulator will need to be installed for the water heaters in the 1″ line. We will discuss more at Tuesday construction meeting.

10/27/04 – PHB reported today that boiler breechings are welded in. One louver is installed, the other will be installed shortly, the blank off plate on the other pipe space will be installed next. The new water softener shipped 10/26. Then unit heater should be shipped very soon. Controls are supposed to be ready to go but TE and PHB will confirm. TE would like to meet on Thursday morning at the site and call Cinergy for the gas pressure issues and begin the resolution process.

10/21/04 – Door will be installed on Monday. Stoermer Anderson will be onsite on Friday. Automation will be on site on Friday. Boiler #1 will be in automatic operation on Friday 10/22/04. PHB will have flues welded in by Wednesday 10/27/04. Boiler # 2 & # 3 should be ready to fire for Cinergy test on 10/28 or 10/29. Please make accommodations for everyone to be onsite Thursday 10/28 for this work. Water softener should ship today or tomorrow for installation.

10/18/04 – Stoermer Anderson reported that gas pressure was very low for the boilers at the site. Boilers require 10 inches. Today’s readings showed 7 inches available. TE is investigating how to solve this problem.

10/18/04 – PHB submitted a price of $4106 to reroute the PVC vents from the domestic water heaters. Do you want to do this work or should we keep things as is and let the trades do the work in the future if it becomes absolutely necessary?

10/15/04 – One boiler successfully started today right on schedule. On Monday SA will verify gas pressure in building as it appeared to be low today. SA is on site preparing boiler for startup. Draft control damper is installed. The door to the boiler room will arrive on Wednesday of next week and will be installed on Wed-Thur-Fri. The insulator will be installing 2″ fiberglass on the steam header to match the 2″ piping shoes, the entire breeching will be replaced with new insulation, A&M is on site downloading the new boiler control program this morning, the CR-1 unit comes factory equipped with a low water cutout switch that turns the pumps off on low CR-1 conditions. TE has the DDC system programmed such that it will stay enabled if a feedwater pump goes into alarm so that when the CR-1 unit comes out of alarm the pumps plant will be standing ready to re-engage. Alarms will be issued on the front end documenting the history. The exhaust fan is being installed today. For future maintenance on the new exhaust the facility group will be able to remove the louver and the fan but they will need a lift to access the fan just like the old one. The fan is not removable from the alley (TE doesn’t want vandals being able to mess with the fan)

10/14/04 – Boiler inspector approved ALL boilers for use today. PHB, TE A&M and SA were all on-site for an impromptu construction meeting when we really were meeting to discuss the installation of the sparge tube in the DA tank. Final decision was to move the vent over to the middle 3″ connection and put a vertical then horizontal sparge tube utilizing a double bushing into the open 3″ connection. SA will do much of the work for this and the temperature sensor for the steam control valve and PHB will provide the final connections to these steam heating points. PHB is finalizing piping and B-1 will be started up on friday towards the very end of the day. B-1 will be put into test service on Monday October 18 where 800 can keep an eye on it during its maiden day of work.

10/12/04 – Mike Martini the boiler inspector will be onsite on Thursday at 10:00a.m. ThermalTech, Peck and Stoermer Anderson would also like to be on site that morning about 8:30 a.m. PHB is trying to get all equipment on site and the boilers ready to fire by the 15th as promised.

10/08/04 – Obtained field data on Chilled Water Pumps, in order to get price quote for replacing seals.

09/03/04 – Date Answered: 09/03/04. Question: The Drip Pan Elbows for the boiler relief vents are shown as 3 inch RV’s. The DPL are being supplied as 3.5″ how should we handle? Response: TE instructed PHB that however the DPL’s are delivered to provide a reducer/increaser to 3″ and run the RV’s as 3″ as they are shown on the P&ID. We believe that the 3.5″ might be the OD listing for the 3″ pipe anyways and may be a non-issue.

09/30/04 – Control equipment is on site except for unit heater control valve. Programming will begin at A&M offices on Oct 3. TE has asked contractors to have one boiler online Oct 8 if possible.

09/29/04 – TTC stopped down and told PHB that they can reuse the existing cond header up to the union, TE prefers that the new boiler stack connections be round to round with a saddle at the breeching connection (remove the long skinny rectangular existing connections, TE wants PHB to provide an allowance price to reroute the domestic water heater PVC exhaust vents to the same window outlet as the gas vents are using (TE will provide a sketch to PHB to price), Boiler #2 should be located to miss the existing pit lid, the EOM traps are correctly sized at 3/4″ so bushings from 2″ to 3/4″ are acceptable, one of the intake louvers appears to be 59″ x 59″ the other window will have the EF in it and the vent lines: PHB can build a painted steel or aluminum blank off plate to surround the vents in the remainder of the window – it has to be water tight to prevent rainwater from entering the space (TE to review CA req’s on remaining louver), TE and PHB reviewed CR-1 inlets and agreed where everything goes, Boiler #4 breeching should just be capped as-is, TE told Tony Martini that existing plant is supplied with combustion air without aid of mechanical combustion air: if 4th boiler is ever added, the County can add a forced combustion air fan, increase existing louver size, or provide another opening in the outside air wall (the existing plant CA req’s are sized on 3 boilers connected load per code req’s)

09/24/04 – Glenwood submitted additional cost based on TE request for strobe relocation and alarm conduit relocation of $214.00 and $427.00. Approved by AKM via phone call.

09/23/04 – Boiler are in the rigger’s yard. Peck will deliver them to 800 on October 2. County, please make appropriations to allow PHB in the building for this work.

09/13/04 – Weekend work was exceptional and the large steel tanks was also removed from the premises.

09/08/04 – PHB Advised that first boiler will be delivered and existing boilers will be demolished on 09/11/04 at 7:00 a.m.

09/03/04 – Control submittals have been received.

09/01/04 – TE approved allowances 01, 02 for PHB shown below.

09/01/04 – TE sent a letter to city explaining boiler plant sizing. We sent email with attachment showing fax.

08/31/04 – PHB has submitted door and louver equipment for approval on 08/24/04 – Returned approved. EF needs to be 203V/3Ph version when ordered.

08/27/04 – TE is trying to get final pricing and ok on water softener package for project from PHB and SA.

08/25/04 – Boiler removal and installation scheduled for September 11, 2004.

08/25/04 – CIS will be removing asbestos in the interim.

08/25/04 – Quick ship boiler scheduled to ship on August 31.

08/25/04 – Condensate Water Return system also scheduled to ship on August 31.

08/25/04 – New door has been cut in with plywood secure doors installed.

08/25/04 – Some prefabricated piping has been deposited on site.

08/25/04 | Date Answered: 08/27/04 Question: A conduit shown to be demolished contains boiler and chiller alarm wiring. Should this be kept and maintained?
Response: TE will confirm with John Noyes next week. The alarms provide remote alarming up to the night desk security shack where the on duty security officer is instructed to call John Noyes. The new control system has been installed with alarm paging service which provides the same functionality so this may not be required at all.

08/12/04 | Date Answered: 08/13/04 – Question: For the spool drawings, what is the most important factor for boiler dimensions.Response: TE stated that the boilers are to aligned under the breeching and centered between the columns but the contractor has some leeway with final and exact boiler placements

by TEI TEI

800 BROADWAY Vacuum Pump Replacement

05/22/09 – We have turned over all the closeout paperwork from Process on this job to Tony. Joe signed the change order at Tony’s office. We created the as-builts and gave them to Bert in CD format with the original marked up drawing. TTC delivered to Bert’s mailbox on 05/18/09.

02/02/09 – TTC inquired about approving this change order again with Darnell & Tony.

01/05/09 – Process is not budging on the a negotiated settlement for their change order request. TTC talked to Joe Doxsey and Tony Matre and could not arbitrate a settlement. TTC asked Joe to call Tony directly to try to convince him of the extra monies. Tony stated the extra valve cost were fine. The sticking point is the extra labor and material mark-up from last year to this year. TTC mailed in payapps 4 & 5 for this proejct which leaves $1000 for closeout plus the change orders to resolve.

12/18/08 – TTC and Process have sent change order payment request to Tony. TTC requested final payment except retainage from Process.

11/25/08 – ThermalTech sent permit drawings to Ginter. Punchlist to Process Const and Change Order from Process to the County.

11/03/08 – Punchlist is completed. Pumps seem to be working fine when all four condensate pumps are left on to handle the building load. TTC has inquired about operation several times with bldg personnel but no response after visit.

10/16/08 – Vacuum pump testing is scheduled for 10/28. Rod will attend for TEI and do a final punchlist.

08/22/08 – TTC talked to Process about project completion which is about 98% complete. TEI asked that O&M’s include warranty letter, equipment submittals and contact information. Commissioning of system needs to happen after the unit is started up in October. Controls should be complete and main pumps should be working. TEI will want to see automatic backup of the pumps and vacuum pumps operate. Each lead pump should work first then each lag pump should take over if there is a failure on the main pumps. Electrical and Controls still have a little work to finish before TE punchlist. Process is submitting a new pay application for most of remaining budget.

08/05/08 – TTC visited the site today to do a pre punchlist check over. The project looks excellent. Piping is nearly complete. Insulation next and then electric and controls. The project should be a couple of weeks from completion. We are investigating the need for additional disconnects.

07/30/08 – Equipment is installed. Insulation and electrical will be installed over next two weeks. Contractor expects to be complete within one month. Process has experienced material cost increases and pay period increases and will be submitting a change order for our review. There is a question about physical disconnects for condensate pumps. EC says we need them. Mechanical schedule says they are provided with the equipment.

07/09/08 – New pipe routing and equipment locations finalized. Housekeeping pad size selected by owner and coordinated with contractor.

Equipment Ordered

Layout changed based on actual footprint of Skidmore equip.

by TEI TEI

800 – Polygonal Pavilion Roof Repair

03-30/12 – THP Reviewed findings with Bert Watts and work complete. THP recommends completing repairs with coating. THP to issue a summary letter.

03/30/12 – THP Reviewed findings with Bert Watts and work complete. THP recommends completing repairs with coating. THP to issue a summary letter.
03/28/12 – THP Reviewed patches during gutter repair project and some patches are cracked. THP to review condition with Bert Wattts.

SSRG completed test patches on west and south elevation. Patches to be reviewed late spring by THP.

by TEI TEI

800 – Gutter Study

03/29/12 – SSRG completed test excavation with SSRG. Drain line not leaking. THP to review findings with Bert Watts. THP with support from SSRG plans to excavate and complete a test repair of gutter at drain line. THP to coordinate work with Darnell.

03/30/12 – THP Reviewed findings with Bert Watts. Water tests did not indicate the gutter or drain line was leaking; but transitions between the gutter and roof are leaking. In conjunction with the roof repairs, additional flashing repairs will be necessary during future projects.

by TEI TEI

800 12th Floor Courtroom Renovations

Project Lead: GBBN

07/11/12 – Bert Watts called and requested GBBN call Tim Moorman to discuss kick-off meeting. GBBN speaks to Tim and discusses project.

07/13/12- GBBN visits the site for the kick-off meeting. GBBN contacts Facilities to discuss original shower reportedly removed from Judge’s Chambers. No one has any knowledge of existing shower on this floor.

07/13/12 – Erica Farris calls requesting another visit to discuss more scope to project. Adding Prosecutor’s Waiting Area renovations or moving of existing Courtroom to other side of floor.

07/13/12 – GBBN verifies existing Facilities AutoCAD files are not fully accurate of existing conditions nor is millwork or furniture shown on these plans.

07/16/12 – GBBN performs field verification of existing spaces and updates AutoCAD drawing files.

07/18/12 – GBBN forwards SOSS to Facilities Department. HCFD emphasizes no work to be performed before SOSS approved.

07/18/12 – Erica Farris requests GBBN to visit site again to discuss additional scope items.

07/18/12 – Erica Farris requests GBBN to forward field verified drawings to Office Furniture Source.

07/19/12 – GBBN forwards updated AutoCAD drawing file via email.

07/23/12 – GBBN visits site again at request of Tim Moorman. Update of project given to Tim upon return from vacation. GBBN asked again about the option of moving the Courtroom to other side of floor. Tim mentioned this option was never discussed previously.

07/30/12 – Erica Farris requests update on design initiation. GBBN reports SOSS has not been approved yet.

07/31/12 – SOSS approved and sent via email from Bert Watts.

08/02/12 – GBBN sends updated AutoCAD files of Office Furniture Source.

08/02/12 – GBBN is contacted by Erica Farris to discuss changes to desgin to add an access corridor from Judge’s Chamber to the Courtroom and changes to the Prosecutor’s Waiting Room. GBBN responds that day.

08/03/12 – ThermalTech sends drawings of the existing 12th Floor mechanical system and description of operation.

08/03/12 – Erica Farris notifies GBBN that Office Furniture Source has suggested subdividing the Prosecutor’s Waiting Room into three space. GBBN is asked to investigate the possibilities. GBBN responds about code issues regarding occupant loads, egress, fresh-air for occupant load for conference rooms and waiting rooms.

08/13/12 – GBBN finalizes Options #1 through #5 along with Pros and Cons of each and an Estimate of Probable Cost for each.

08/17/12 – GBBN meets with Erica Farris, Judge Hunter, Case Manager and Sheriff’s Deputy to review options and discuss other programming issues. Options are reviewed and critiqued. GBBN is instructed to modify Option in Courtroom to include Deputy location and Additional Attorney Seating.

08/23/12 – GBBN sends updated Option #6 showing input from previous site meeting.

08/24/12 – JAVS contacted by GBBN about relocation of existing control box and Opinion of Probable Cost via email on website.

08/26/12 – Bert Watts is sent all Design Options and Opinions of Probable Cost.

08/27/12 – JAVS contacts GBBN about inquiry into relocation of existing control box and Opinion of Probable Cost. Future site visit may be coordinated in near future. Bert Watts indicates Judge Hunter’s Office is now requesting painting of Courtroom and adjacent spaces, refinishing of existing hardwood floors and carpet/base be added to scope. GBBN forwards pricing information for painting to Bert Watts.

08/29/12 – Erica Farris requests GBBN to visit the site again to discuss additional changes to the Option #6. GBBN visits site to meet with Erica Farris. The design is modified and resent via email that afternoon.

09/07/12 – Bert Watts and GBBN meet with Judge Hunter and Erica Farris to finalize latest schemes and options for the Courtroom and Attorney’s Waiting Areas. Discussions held to examine how to proceed from this point forward.

by TEI TEI

800 Electrical Power Failure Emergency Response

Project Lead: TEI

01/24/14 – Due to sprinkler line freeze and break on the second floor, water infiltrated the electrical gear in the Sub Basement of the the building and caused a fault, which then left the building without utility power for 2-1/2 days. Hamilton county facilities department, ThermalTech, Duke, IBI, county electricians, and building staff worked collaboratively with each other to restore utility power to the building. The fault left the building with only one Duke feed (The Duke switching cabinet in the sub basement requires that one of the switches be re-built by Duke). The Control Power transformer in the building service was damaged. Jeff Jones from ECI has placed order for replacement CTT.

01/24/14 – TEI submits SOSS to Bert Watts.

02/25/14 – TEI receives approved SOSS and purchase order.

03/26/14 – TEI completes work associated with emergency response.

04/24/14 – Project Complete.

by TEI TEI

800 Electrical Switchgear Repair

Project Lead: TEI

01/24/14 – TEI communicating with CEPower and IBI to determine regulatory regulations that will be required to restore both services to operation (normal and standby service from Duke). TEI will then author and submit SOSS to Bert Watts for approval.

02/25/14 – TEI met with Matt Kennedy of CEPower on 2/20/14 to discuss scope of services that CEPower to provide. CEP and TEI had a conference call on 02/25/14 to discuss project, test plans, etc. TEI to submit SOSS to Bert watts upon receipt of CEP fees (to work as a sub to TEI). Anticipated ship date for CPT from General Electric is 03/13/14 (according to Jeff Jones).

03/26/14 – TEI received late on 03/25/14 proposal from CE Power to do testing. TEI authoring the SOSS for working with IBI, Jeff Jones, and CE Power to restore power to remainder of County owned equipment. TEI communicated with Duke Energy (Jeff Hesse) to advise that the anticipated restore power date from the County switchgear. Jeff Jones of CEI indicated that the replacement control power transformer and relays are at his shop and as soon as PO and contracts are executed, the re-install can be scheduled.

04/22/14 – SOSS submitted to HamCo. Project approved and commencing immediately. Duke repairs have began. TEi to schedule Work on getting the Permit completed sot that the gear can be restored to full pre-fault functionality.

05/28/14 – ThermalTech has met with CEPower and discussed with Jeff Jones the restoration of the switchboard. CEPower to identify length of time that the outage will take and then also give ThermalTech timing of when the testing can occur. Jeff Jones has equipment at his shop and plans is to install all parts that are in non-energized cabinets ahead of outage to keep the outage duration to a minimum. HamCo to identify timing when outage can and/or can not occur.

06/25/14 – The switchgear is restored on the County side. IBI inspection on 06/25/14 (approved – temporary release) and IBI to send release to Duke to get release to power up gear for final testing by CE Power. Duke anticipates to restore their switching and be prepared to get power back on week of June 30, 2014.

07/30/14 – Duke switchgear restored and ThermalTech received notification on 07/14/14 that Duke was ready to re-energize. ThermalTech contacted all of the parties (CEP, ECI, IBI HamCo Duke) and the final testing and restoration of the gear is scheduled for 08/09/14, work beginning at 8:00 AM.

08/25/14 – All associated parties met on site on Saturday 08/09/14. The primary and secondary sources are now connected and the auto-throw over equipment is fully functioning per the original design intent of the gear. IBI inspected the gear and gave a final approval of the restoration of the gear. power to both primary and standby services are now on and the building gear is back to being fully functional.

by TEI TEI

800 Energy Star 2014

Project Lead: TTE

09/23/14 – Terry and Bracken visited the building and took the air, people and light readings for the verification process.

10/23/14 – Final submission paperwork is complete. Ralph signed, ThermalTech is uploading to website.

12/12/14 – Building was certified in November and uploaded to LEED with an 88 score.

by TEI TEI

800 Freezestats

Project Lead: ThermalTech

08/29/12 – ThermalTech visited the site several times to discuss 2013 installation of freezestats by Process Construction. It was determined and recommend by TTE that freezestats on units which have had their non-glycol preheat hot water coils removed do not need freezestats reinstalled. This applies to most of the low rise units. A&M will leave the freezestats in the units and jumper across them so they don’t shut the fans down on morning warmup but will also implement a full morning warmup cycle in cold weather to prevent unnecessary overcooling of the Supply Air stream in these units.

11/28/14 – After investigating the L-71 rooftop unit with the same issue and getting prices from the vendor for a replacement unit or a electric preheat coil option it was determined the same strategy could be used for this unit. The freezestat will be bypassed, this was discussed and agreed on between Terry Cannon, Darnell Edward and Tony Matre. The unit has not frozen in the 14 years we have been working with the County at this building. The RF-1 also does not need to run to prevent the freezestat from tripping, A&M will bypass this freezestat too.

by TEI TEI

800 Lighting Study and Report

Project Lead: ThermalTech

08/29/12 – SOSS approved and PO received by ThermalTech. ThermalTech to work with BW and schedule kickoff to determine study schedule and needs.

09/26/12 – Initial daytime readings (Normal power) completed by ThermalTech on 9/25. Mike Barth reviewing findings and preparing night time night light and emergency light readings schedule. Coordinating with Darnell Edwards. Initial conclusions – most places are within guidelines, however, there are many areas that lighting is low and many other places where lighting levels are high.

10/31/12 – Night light night readings are scheduled to occur between 8:00 PM and 6:00 AM on Thursday night, November 1st, 2012. Mike Barth of ThermalTech coordinated with building manager and Tony Matre for access to building. Upon completion of the survey, the results and report will be written up and sent to Bert Watts for review.

01/30/13 – All light readings have been complied and indicated on drawings. Report is in process and will be completed in next few weeks for Bert Watts to review.

04/24/13 – Review copy of report to be sent to BW by end of April for comments.

03/25/15 – TEI received BW review comments, and printed 3 copies of final report for Bert Watts. Final report delivered to BW on 03/20/15. Project complete and can be archived.

by TEI TEI

800 Roof Repair

Project Lead: THP

06/01/12 – THP submitted SOSS.

09/13/12 – County approves SOSS.

10/04/12 – THP County kick-off meeting complete. DD documents due early November and Final bid documents due mid-December with a budget. Bid period mid-February.

10/31/12 – THP working on DD documents.

01/15/13 – THP Issued DD documents for review.

02/08/13 – County returned project manual comment.

02/19/13 – THP and County had a design review meeting.

05/17/13 – THP issued 95% Review documents to the County.

06/27/13 – Project update. Due to weather sensitive work, if project approved work will not occur until next spring. Possible bid project late 2012.

10/31/13 – Project on hold.

01/30/14 – Project on hold.

04/24/14 – Project on hold.

06/26/14 – Project on hold.

07/31/14 – Project on hold.

08/27/14 – Project on hold.

09/25/14 – Overall project on hold.

01/29/15 – Overall project on hold.

02/26/15 – Overall project on hold.

03/26/15 – Overall project on hold

04/01/15 – THP requested to assess current roof condition and provide a summary report letter.

04/17/15 – THP published roof assessment letter with updated opinion of probable cost.

by TEI TEI

800 Domestic Relations Renovation

Project Lead: GBBN

06/23/15 – County requested GBBN attend a kick-off Meeting at 800 Broadway on Tuesday, June 30th at 2:00 pm.

06/30/15 – Kick-off Meeting held at Domestic Relations area. Scope drawings prepared by the users are distributed.

07/02/15 – Original SOSS submitted to Facilities Department.

07/15/15 – John Nester contacts GBBN about revising SOSS without walk-through permitting services.

07/16/15 – GBBN submits revised SOSS to Facilities Department.

07/30/15 – Approval came in yesterday. They would like to move in by September. Try to make renovation require no sprinkler work so split the sprinklers with walls.

08/18/15 – TEI confirms no fire suppression sprinkler head modifications are required on this project. Bert Watts finalizes the last scope questions to GBBN.

08/20/15 – GBBN submits combined Building and Mechanical Permit to the City of Cincinnati Permit Center.

08/25/15 – TEI submitted drawings to IBI for approval, pending plan review. TEI reviewed CAGIS page and plan review issued complete (CIC1503111).

09/23/15 – John Nester indicated that the electricians are ready for a rough-in inspection. TEI contacted IBI, and IBI can not perform any inspections until the building permit is issued. TEI is waiting on notification that the building permit has been issued to contact IBI again. HVAC is complete and measured airflows are at or above design CFM’s.

10/28/15 – Rough in inspection completed and passed with IBI. ThermalTech called IBI for the electrical final inspection, scheduled for Friday morning, 10/30/15. ThermalTech to contact IBI Friday between 7:30 and 8:30 for inspection timeframe window.

by TEI TEI

800 Energy Star

Project Lead: TEI

10/27/15 – TTC visited the site yesterday for validation. Everything looks great.

12/08/15 – Energy Star Certification complete.

by TEI TEI

800 6th Fl Courtroom Renovations

Project Lead: GBBN

08/12/13 – Kick-off Meeting with Bert Watts and ThermalTech.

08/16/13- TEI submits break down of scope, reimbursable expenses and hours to GBBN,

08/17/13- County requests new SOSS based on lowered scope. Bert Watts confirms no scope of work involved on 12th Floor. Moving by Owner.

08/19/13- GBBN submits SOSS to Bert Watts.

08/19/13 – County requests break downs for consultant scope and hours for the SOSS to GBBN. ThermalTech is working on the CT strategy and will talk to the HVC chemical treatment contractor.

05/21/14 – GBBN requests full recycling tracking information beyond the paper only sent 5/24/14. Request also made for total waste disposal records from building to determine percentage of waste that is recycled.

05/27/14 – Bert Watts sends Sustainable Purchasing Tracking information to GBBN.

05/29/14 – GBBN receives word from USGBC that the review of the LEED EB+OM application has been received.

06/10/14 – GBBN receives word from USGBC that the review of the LEED EB+OM application is commencing.

07/09/14 – USGBC sends initial review comments to GBBN for review.

07/28/14 – All involved partners attend team meeting to discuss approach for returning responses to the review comments back to USGBC.

09/09/14 – Bert Watts requests estimate for this project to be authored and billed to General Services Contract.

01/28/16 – Project ready to archive.

by TEI TEI

800 First Floor Staff Relocation

Project Lead: GBBN

08/09/15 – County requested GBBN attend a kick-off Meeting at 800 Broadway on Tuesday, August 11th at 2:00 pm.

08/11/15 – Kick-off Meeting held at 1st floor north Office Suite area. Scope drawings prepared by Facilities is distributed to GBBN. Photos taken of north and south side areas and existing systems furniture.

08/12/15 – GBBN notifies Facilities of the existing electrical code implications of disassembling and reassembling the existing workstation cubicles per TEI and IPM’s input.

08/25/15 – Facilities notifies GBBN of the two Archibus Pop-up Project Request numbers for the estimating input.

08/26/15 – GBBN completes both estimates in Archibus and reports to Facilities they are completed.

by TEI TEI

800 Court of Domestic Relations Wireless Study

Project Lead: TEI

09/23/15 – TEI submitted SOSS to Bert Watts, Pending Approval.

09/24/15 – Bert says they might already have some wireless in place so this scope may be altered.

10/28/15 – TEI gets approved SOSS and PO – Bert and TEI trying to schedule Kick off meeting with Shannon Supe et al.

12/07/15 – TEI met with Greg Wandstatt and Todd Coriell to review the requirements from the Tele-Comm department. TEI met with Lisa Gorrasi (Court Administrator – Domestic Relations) on 12/08/15 to walk the space and to determine where in the building the existing WAP are located and where the areas of low/no coverage are. TEI to coordinate final recommendations with Facilities and with Shannon Supe.

01/26/16 – TEI issues copy of study to Bert Watts for review. Bert Watts approves study and asks for cost to implement the recommendations. TEI to develop cost and send to Bert Watts.

02/23/16 – TEI issues final report to Bert Watts for distribution.

by TEI TEI

800 – Building Envelope Assessment

Project Lead: THP

06/13/16 – Stone debris was discovered on the east elevation. THP requested to complete an emergency stone review then a more formal building envelope assessment. ZSR to provide contractor support.

06/17/16 – THP submitted an SOSS.

06/21/16 – County approved SOSS.

06/22/16-06/28/16 – Emergency inspections of the façade via aerial platform and rappelling.

06/22/16 – THP started five day aerial survey of the base of the building and rappelling survey of the tower. THP provide details to stabilize (4) critical stone conditions on the base of the building.

06/27/16 – THP issued sketch of locations of stones to receive supplemental anchors. Waiting on some additional scans from archives.

07/25/16 – THP published a worst case stone and angle repair budget for the County to consider during short term planning.

07/28/16 – Final report anticipated to be complete mid-September.

01/26/17 – THP published final report.

02/03/17 – THP published report to County.

02/17/17 – THP updated cost estimate and published revision to final report.

by TEI TEI

800 Flat Roof Leak Assessment

Project Lead: THP

02/16/17 – Hamilton County requested SOSS.

03/10/17 – Met with Hamilton County to review scope and field conditions.

03/23/17 – THP provided Hamilton County with a SOSS.

04/06/17 – SOSS approved.

04/27/17 – Thermal scan rained out earlier in April. Next scheduled scan is May 2nd (weather permitting).

05/23/17 – THP completed thermal scan of roof.

05/25/17 – THP visited site to meet with County to observe and document areas of flat roofs where water was trapped behind membrane at south-facing parapet wall.

06/27/17 – County and THP to review active leak areas.

06/29/17 – THP working on report.

07/27/17 – THP working on report.

08/31/17 – THP published report for review by County.

by TEI TEI

800 Adult Probation Sub-Basement Project

Project Lead: ThermalTech

02/22/17 – TEI and Facilities met with Assistant Chief Probation Officer to scope out the project for the renovation of existing paper storage room. TEI to develop SOSS and submit to BW by 03/03/17.

03/20/17 – Met onsite to do more fieldwork. We found plumbing, HVAC and electrical locations for this work. Our belief after this field trip is that we can make this job work in this room.

03/22/17 – Project in design – fast track project to meet client deadlines.

04/12/17 – THP ThermalTech requested structural review of the proposed unit on the raised floor.

04/26/16 – Permits are in for electric and mechanical. Queen City Mechanical is submitting a permit for the plumbing but it should also be in by now. Mechanical bids have been submitted and forwarded to Bert Watts.

04/26/17 – THP published findings that design capacity of the raised flooring would support the listed weight of the AU480.

04/27/17 – Electrical Plan review approved. ThermalTech to pick up permit and drop off with facilities prior to monthly partner meeting.

05/22/17 – Electrical permit picked up and dropped off with facilities department.

06/08/16 – Feldkamp asked if they can route the new ventilation line form the room duct in flex. TTC reluctantly agreed but only if it can bring in the 50cfm required on the drawings.

06/22/17 – Electrical Installation complete, and the project has been inspected and final IBI approval received.

06/23/17 – Talked to Feldkamp and HVAC is complete, they were unclear about submitting for their own electrical IBI permit just for the unit but they are taking care of that.

07/25/17 – TEI provided Feldkamp revised drawing for IBI approval. TEI Checked EZ Track (Cagis) at 3:00 PM on 07/25 and no change in status (last update on Cagis was 06/26).

07/27/17- Electrical inspection complete, plumbing approved waiting for startup, mechanical permit approved, electrical HVAC permit awaiting approval. A backflow preventer may be added after startup of the new equipment.

08/31/17 – HVAC permit has deficiency concern for refrigerant in the boiler room ancillary room. TTC to check with Feldkamp.

09/25/17 – Called the City Permit department about comment on condenser being in mechanical room. I don’t understand this comment as R-410a has no flame propagation potential and its less than 5 lbs so there should be no issue with permitting, plus there is no OA for AHU in this space. I sent all this info to Feldkamp HVAC as well hoping to resolve this issue.

10/26/17 – Pat says this will be wrapped up by the week’s end.

11/28/17 – Feldkamp says that permit is cleared up with Reggie at the City and they will get this job closed out. TTC contacted them today to push that along. Reggie has clossed and approved the project.

by TEI TEI

800 – 2015 Roof Repair – Cleaning

Project Lead: THP
04/13/15 – County requested SOSS.
04/28/15 – THP issued SOSS.
07/31/15 – County issued approved SOSS.
08/26/15 – Project update: THP anticipates issuing 90% Owner review drawings by 8/31.
09/21/15 – THP issued DD documents.
09/22/15 – County provided comments to DD documents.
11/04/15 – Prebid meeting occurred.
11/13/15 – Addendum No. 1 issued.
11/14/15 – County – three bids from Midwest Maintenance, SSRG and ZSR.
11/30/24 – THP issued a letter recommending ZSR, based on the low bid and follow up conversation.
11/20/15 – THP issued draft AIA contract.
12/08/15 – County provided edits to contract for THP to update.
01/28/16 – Project summary. ZSR – final AIA contract, signed it and sent to the County for execution. ZSR submitted tentative submittals.
01/27/16 – County commissioners approved ZSR contract.
02/10/16 – THP returned submittals.
02/25/16 – ZSR finalizing construction schedule. THP trying to work with SHPO to coordinate anticipated work.
04/28/16 – Project update: ZSR plans to start work later part of May.
05/05/16 – THP makes contact with SHPO. They direct THP to regional manager. THP emails John Sandor of “Technical Preservation Services” for project review.
05/10/16 – Preconstruction Meeting Held On-Site
05/16/16 – Project began with erection of scaffolding and sidewalk protection
05/23/16 – THP makes site visit to review scope with superintendent
05/24/16 – Weekly meeting Held On Site
05/24/16 – THP receives contact back from John Sandor. THP to issue email to John S. summarizing conversation. John S. will issue a letter on the project with his official recommendations.
05/25/16 – Scaffolding complete
05/31/16 – ZSR to start construction.
06/21/16 – Construction progress meeting
06/28/16 – Limestone demolition is almost complete.
06/29/16 – Revised limestone anchorage detail issued by THP for pricing.
07/05/16 – Met project team for progress meeting and to document progress.
07/08/16 – THP developed AIA change orders 1 & 2 and forwarded to ZSR.
07/08/16 – THP issued 7/7 progress meeting notes and field report.
07/12/16 – Developed kerf edge repair detail and forwarded to ZSR.
07/12/16 – ZSR not on site, waiting for materials to arrive.
07/19/16 – Met project team for progress meeting and to document progress.
07/27/16 – Bert confirmed ZSR has – payment for app. #1.
07/26/16 – THP issued 7/19 meeting notes and field report.
07/27/16 – Met ZSR at site to review progress, and stone profile/alignment issues.
07/28/16 – Demolition of stone panels at the bottom is complete. Stone ordered with 8 week delivery anticipated. Crews removing the metal caps.
07/28/16 – THP visited site and met with Darnell to review and document water infiltration.
07/29/16 – THP issued field report 07 from 7/27 & 28 site visits.
08/02/16 – Met project team for progress meeting and to document progress.
08/03/16 – THP researched specified patch and consolidation treatment materials.
08/08/16 – THP issued 8/2 meeting notes and field report.
08/08/16 – THP researched specified water repellent.
08/10/16 – THP emailed ZSR following up on outstanding items from field report 08.
08/11/16 – THP visited site to document progress, and issued field report 09.
08/16/16 – THP emailed ZSR dimensioned sketch of typical stone panel showing fabrication details to facilitate installation.
08/17/16 – THP visited site and met with Darnell to review and document water infiltration.
08/18/16 – THP issued field report 10 from 8/17 site visit.
08/23/16 – THP visited site to observe and document roof progress.
08/25/16 – THP issued field report 11 from 8/23 site visit.
08/26/16 – THP inspected stone patches at East roof face while rappelling to assess tower facades.
08/30/16 – Progress meeting and THP site visit to document progress.
09/06/16 – THP issued 8/30 meeting minutes and field report 12.
09/09/16 – THP made two visits to site to document progress and review stone profile with ZSR, and issued field report 13.
09/13/16 – Progress meeting and THP site visit to document progress. THP coordinated revised installation detail for stones due to field conditions.
09/14/16 – THP issued 9/13 meeting minutes and field report 13.
09/27/16 – Progress meeting and THP site visit to document progress, and issued meeting minutes and field report 13.
10/03/16 – THP finalized approval of ZSR’s Applications for Payment 2, 3, and 4.
10/06/16 – THP – and reviewed Stone Center shop drawings.
10/07/16 – THP met with ZSR to review Stone Center shop drawing to identify what changes needed to be made.
10/11/16 – Progress meeting, and THP site visit to observe and document progress.
10/13/16 – THP – and reviewed revised Stone Center shop drawing, and returned to ZSR to be revised and resubmitted.
10/18/16 – THP issued 10/11 meeting minutes and field report 16.
10/20/16 – THP – and reviewed revised Stone Center shop drawing, and returned to ZSR to be revised and resubmitted.
10/21/16 – THP – and reviewed final Stone Center shop drawing, and returned to ZSR with exceptions noted.
10/25/16 – THP – and reviewed ZSR’s Application for Payment #5, and forwarded to Hamilton County.
10/26/16 – THP site visit to observe and document progress.
11/01/16 – THP issued field report 17.
11/01/16 – Progress meeting, and THP site visit to observe and document progress.
11/01/16 – THP issued 11/1 meeting minutes and field report 18.
11/02/16 – THP finalized approval of ZSR’s Applications for Payment 2-5 and delivered to Hamilton County.
11/15/16 – Progress meeting, and THP site visit to observe and document progress.
11/17/16 – THP issued 11/15 meeting notes and field report 19.
11/18/16 – THP site visit to observe preformed sealant strip mockups, and issued field report 20.
11/29/16 – Progress meeting, and THP site visit to observe and document progress.
11/30/16 – THP issued 11/29 meeting minutes and field report 21
12/03/16 – 12/04/16 – ZSR installed new limestone roofing.
12/06/16 – THP reviewed ZSR’s proposals for change orders 3 & 4 and replied with revisions required per contract.
12/09/16 – THP issued field reports 21 & 22, and delivered Application for Payment 6 and Change Orders 3 & 4 to Hamilton County.
12/27/16 – Progress meeting, and THP site visit to observe and document progress. THP issued 12/27 meeting notes and field report 24.
12/28/16 – THP provided ZSR information about stone cleaning products to be used for developing test areas.
01/04/17 – ZSR finished installation of snow guards.
01/16/17 – American Scaffolding began removing scaffolding.
01/23/17 – ZSR site visit to document removal of scaffolding and progress of preformed sealant strip installation.
01/23/17 – ZSR site visit to document removal of scaffolding and progress of preformed sealant strip installation, and issued field report 25.
02/22/17 – Project on hold due to weather conditions.
03/01/17 – THP visited site to document leaks at 17th and 18th floor mechanical spaces.
03/02/17 – THP issued field report #26 with plan diagrams identifying leak locations.
04/11/17 – THP visited the site during a rain event to document wall and roof leaks. Findings and recommendations emailed to County.
04/17/17 – THP and ZSR discussed plan to test leak areas and testing of stone cleaning.
05/12/17 – THP visited site to observe and document progress of work and test sample cleaning areas.
05/12/17 – THP issued field report #27.
05/17/17 – THP met ZSR and Prosoco product rep at site to discuss options for additional stone cleaning products.
05/19/17 – THP emailed additional stone cleaning products to ZSR for test sampling.
05/25/17 – ZSR completed premold and sealant installation. Material ordered for cleaning samples. Cleaning probably occur next week (weather permitting).
06/08/17 – THP visited site to coordinate with ZSR application of limestone cleaning product for testing, and document progress of roof repairs.
06/08/17 – THP issued field report #28 documenting cleaning test areas, methods, and products.
06/09/17 – THP issued field report #29 documenting progress and identifying additional work items.
06/13/17 – THP visited site to sound stone patches and document results of stone cleaning tests.
06/27/17 – Meeting at site with Hamilton County, ZSR, and THP to review results of stone cleaning tests. THP recommended two samples for additional testing and pricing.
07/27/17 – County approved tentative design and construction budget for cleaning change order work. ZSR and THP exploring staining patches vs. cleaning more effective option.
08/31/17 – As part of the cleaning repair project, ZSR complete the general construction cleaning. Refer to F98.

 

by GBBN GBBN

800 – Juvenile Court and Domestic Court Relocation Study

06/06/17 – GBBN contacted by Facilities about authoring SOSS.
06/07/17 – Hamilton County Facilities requests some modifications to the SOSS including using #F052.
06/08/17 – Revised SOSS submitted to Facilities.
06/27/17 – Facilities sends approved SOSS to GBBN.
06/28/17 – Facilities and GBBN begin discussions for a Kick-off Meeting date.
08/02/17 – Meeting with Facilities and GBBN to discuss scope and approach for moving ahead with project.
08/05/17 – GBBN holds internal review meeting before meeting for presentation to Owner.
08/17/17 – Meetings at 800 Broadway with both Juvenile Court and Domestic Relations Court. Tour of Domestic Court Offices was conducted after meeting. Domestic Relations presented Programming List for GBBN’s use.
08/25/17 – GBBN internally produces meeting minutes and circulates for review.
09/22/17 – GBBN holds internal review meeting before meeting for presentation to Owner.
10/10/17 – Final presentation review set distributed in office for associated personnel to review and comment.
10/18/17 – GBBN communicates with Facilities and Administration related to timing for meeting for final presentation. A final Onwer review draft is sent to the Owner. Ralph Linne responded with corrected financial information.
10/20/17 – Corrected draft with updated financial information is sent by GBBN to all parties involved.

by TEI TEI

800 – EMU Relocation Study

Project Lead: GBBN
03/28/17 – GBBN contacted by Facilities about setting up Kick-off Meeting with Sheriff Department.
03/31/17 – Hamilton County Facilities and GBBN meet with Sheriff to discuss project goals. Letter from Sheriff requesting project presented to GBBN. Facilities requested SOSS from GBBN.
03/10/17 – Met with Hamilton County to review scope and field conditions.
03/23/17 – THP provided Hamilton County with a SOSS.
04/05/17 – GBBN visits site for field verification of measurements and to take photographs.
04/06/17 – GBBN corresponds with Facilities for final questions related to scope of work. GBBN authors SOSS and delivers to Facilities.
04/14/17 – SOSS approved.
05/04/17 – GBBN revisits site for additional field verification, measurements and to take more detailed photographs.
05/05/17 – GBBN interviews ThermalTech for costing details on Electrical and Mechanical issues.
05/11/17 – GBBN sends six (6) optional designs to Facilities for review.
05/12/17 – Facilities selects two final options for estimating. One with the existing demising wall to remain and one without the existing wall present.
05/22/17 – GBBN completes preliminary Opinions of Probable Costs for each of the two final options.
05/23/17 – Facilities indicates only one option, with the existing demising wall removed, is required for estimating.
06/14/17 – GBBN, Facilities and EMU meeting to discuss the version of the schematic design selected and estimated. Minor modifications were requested for the final option.
07/18/17 – Facilities indicated this project has been terminated. The move may result in relocating to the Third Floor Mezzanine instead.
07/24/17 – TEI discussed project with County and GBBN and is doing an energy study for the 24/7 part at the very least. The Energy Impact Study has been submitted for this project.
12/19/17 – TTC adds a comment.

by TEI TEI

800 – 2015 Roof Repair

Project Lead: THP
04/13/15 – County requested SOSS.
04/28/15 – THP issued SOSS.
07/31/15 – County issued approved SOSS.
08/26/15 – Project update: THP anticipates issuing 90% Owner review drawings by 8/31.
09/21/15 – THP issued DD documents.
09/22/15 – County provided comments to DD documents.
11/04/15 – Prebid meeting occurred.
11/13/15 – Addendum No. 1 issued.
11/14/15 – County – three bids from Midwest Maintenance, SSRG and ZSR.
11/30/24 – THP issued a letter recommending ZSR, based on the low bid and follow up conversation.
11/20/15 – THP issued draft AIA contract.
12/08/15 – County provided edits to contract for THP to update.
01/28/16 – Project summary. ZSR – final AIA contract, signed it and sent to the County for execution. ZSR submitted tentative submittals.
01/27/16 – County commissioners approved ZSR contract.
02/10/16 – THP returned submittals.
02/25/16 – ZSR finalizing construction schedule. THP trying to work with SHPO to coordinate anticipated work.
04/28/16 – Project update: ZSR plans to start work later part of May.
05/05/16 – THP makes contact with SHPO. They direct THP to regional manager. THP emails John Sandor of “Technical Preservation Services” for project review.
05/10/16 – Preconstruction Meeting Held On-Site
05/16/16 – Project began with erection of scaffolding and sidewalk protection
05/23/16 – THP makes site visit to review scope with superintendent
05/24/16 – Weekly meeting Held On Site
05/24/16 – THP receives contact back from John Sandor. THP to issue email to John S. summarizing conversation. John S. will issue a letter on the project with his official recommendations.
05/25/16 – Scaffolding complete
05/31/16 – ZSR to start construction.
06/21/16 – Construction progress meeting
06/28/16 – Limestone demolition is almost complete.
06/29/16 – Revised limestone anchorage detail issued by THP for pricing.
07/05/16 – Met project team for progress meeting and to document progress.
07/08/16 – THP developed AIA change orders 1 & 2 and forwarded to ZSR.
07/08/16 – THP issued 7/7 progress meeting notes and field report.
07/12/16 – Developed kerf edge repair detail and forwarded to ZSR.
07/12/16 – ZSR not on site, waiting for materials to arrive.
07/19/16 – Met project team for progress meeting and to document progress.
07/27/16 – Bert confirmed ZSR has – payment for app. #1.
07/26/16 – THP issued 7/19 meeting notes and field report.
07/27/16 – Met ZSR at site to review progress, and stone profile/alignment issues.
07/28/16 – Demolition of stone panels at the bottom is complete. Stone ordered with 8 week delivery anticipated. Crews removing the metal caps.
07/28/16 – THP visited site and met with Darnell to review and document water infiltration.
07/29/16 – THP issued field report 07 from 7/27 & 28 site visits.
08/02/16 – Met project team for progress meeting and to document progress.
08/03/16 – THP researched specified patch and consolidation treatment materials.
08/08/16 – THP issued 8/2 meeting notes and field report.
08/08/16 – THP researched specified water repellent.
08/10/16 – THP emailed ZSR following up on outstanding items from field report 08.
08/11/16 – THP visited site to document progress, and issued field report 09.
08/16/16 – THP emailed ZSR dimensioned sketch of typical stone panel showing fabrication details to facilitate installation.
08/17/16 – THP visited site and met with Darnell to review and document water infiltration.
08/18/16 – THP issued field report 10 from 8/17 site visit.
08/23/16 – THP visited site to observe and document roof progress.
08/25/16 – THP issued field report 11 from 8/23 site visit.
08/26/16 – THP inspected stone patches at East roof face while rappelling to assess tower facades.
08/30/16 – Progress meeting and THP site visit to document progress.
09/06/16 – THP issued 8/30 meeting minutes and field report 12.
09/09/16 – THP made two visits to site to document progress and review stone profile with ZSR, and issued field report 13.
09/13/16 – Progress meeting and THP site visit to document progress. THP coordinated revised installation detail for stones due to field conditions.
09/14/16 – THP issued 9/13 meeting minutes and field report 13.
09/27/16 – Progress meeting and THP site visit to document progress, and issued meeting minutes and field report 13.
10/03/16 – THP finalized approval of ZSR’s Applications for Payment 2, 3, and 4.
10/06/16 – THP – and reviewed Stone Center shop drawings.
10/07/16 – THP met with ZSR to review Stone Center shop drawing to identify what changes needed to be made.
10/11/16 – Progress meeting, and THP site visit to observe and document progress.
10/13/16 – THP – and reviewed revised Stone Center shop drawing, and returned to ZSR to be revised and resubmitted.
10/18/16 – THP issued 10/11 meeting minutes and field report 16.
10/20/16 – THP – and reviewed revised Stone Center shop drawing, and returned to ZSR to be revised and resubmitted.
10/21/16 – THP – and reviewed final Stone Center shop drawing, and returned to ZSR with exceptions noted.
10/25/16 – THP – and reviewed ZSR’s Application for Payment #5, and forwarded to Hamilton County.
10/26/16 – THP site visit to observe and document progress.
11/01/16 – THP issued field report 17.
11/01/16 – Progress meeting, and THP site visit to observe and document progress.
11/01/16 – THP issued 11/1 meeting minutes and field report 18.
11/02/16 – THP finalized approval of ZSR’s Applications for Payment 2-5 and delivered to Hamilton County.
11/15/16 – Progress meeting, and THP site visit to observe and document progress.
11/17/16 – THP issued 11/15 meeting notes and field report 19.
11/18/16 – THP site visit to observe preformed sealant strip mockups, and issued field report 20.
11/29/16 – Progress meeting, and THP site visit to observe and document progress.
11/30/16 – THP issued 11/29 meeting minutes and field report 21
12/03/16 – 12/04/16 – ZSR installed new limestone roofing.
12/06/16 – THP reviewed ZSR’s proposals for change orders 3 & 4 and replied with revisions required per contract.
12/09/16 – THP issued field reports 21 & 22, and delivered Application for Payment 6 and Change Orders 3 & 4 to Hamilton County.
12/27/16 – Progress meeting, and THP site visit to observe and document progress. THP issued 12/27 meeting notes and field report 24.
12/28/16 – THP provided ZSR information about stone cleaning products to be used for developing test areas.
01/04/17 – ZSR finished installation of snow guards.
01/16/17 – American Scaffolding began removing scaffolding.
01/23/17 – ZSR site visit to document removal of scaffolding and progress of preformed sealant strip installation.
01/23/17 – ZSR site visit to document removal of scaffolding and progress of preformed sealant strip installation, and issued field report 25.
02/22/17 – Project on hold due to weather conditions.
03/01/17 – THP visited site to document leaks at 17th and 18th floor mechanical spaces.
03/02/17 – THP issued field report #26 with plan diagrams identifying leak locations.
04/11/17 – THP visited the site during a rain event to document wall and roof leaks. Findings and recommendations emailed to County.
04/17/17 – THP and ZSR discussed plan to test leak areas and testing of stone cleaning.
05/12/17 – THP visited site to observe and document progress of work and test sample cleaning areas.
05/12/17 – THP issued field report #27.
05/17/17 – THP met ZSR and Prosoco product rep at site to discuss options for additional stone cleaning products.
05/19/17 – THP emailed additional stone cleaning products to ZSR for test sampling.
05/25/17 – ZSR completed premold and sealant installation. Material ordered for cleaning samples. Cleaning probably occur next week (weather permitting).
06/08/17 – THP visited site to coordinate with ZSR application of limestone cleaning product for testing, and document progress of roof repairs.
06/08/17 – THP issued field report #28 documenting cleaning test areas, methods, and products.
06/09/17 – THP issued field report #29 documenting progress and identifying additional work items.
06/13/17 – THP visited site to sound stone patches and document results of stone cleaning tests.
06/27/17 – Meeting at site with Hamilton County, ZSR, and THP to review results of stone cleaning tests. THP recommended two samples for additional testing and pricing.
07/27/17 – County approved tentative design and construction budget for cleaning change order work. ZSR and THP exploring staining patches vs. cleaning more effective option.
08/31/17 – As part of the cleaning repair project, ZSR complete the general construction cleaning. Refer to F98.